### v2

# The metric theory of small gaps for a sequence of real numbers

Abstract.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive integers. The metric theory of minimal gaps for the sequence $\{\alpha a_{n}\text{ mod }1,1\leq n\leq N\}$ as $N\to\infty$ was initiated by Rudnick [Rudnick], who established that the minimal gap admits an asymptotic upper bound expressible in terms of the additive energy of $\{a_{1},\ldots,a_{N}\}$ for almost every $\alpha$. Later, Aistleitner, El-Baz, and Munsch [ABM] demonstrated that the metric theory of minimal gaps for such sequences is not governed by the additive energy, but by the cardinality of the difference set of $\{a_{1},\ldots,a_{N}\}$. They established a sharp convergence test for the typical asymptotic order of the minimal gap and proved general upper and lower bounds that are readily applicable. A key element of their proof relies on the resolution of the Duffin–Schaeffer conjecture by Koukoulopoulos and Maynard. In this article, we generalise several results from [ABM] on integer sequences to the case of real sequences. Although an upper bound for $\delta_{\min}^{\alpha}(N)$ remains elusive, we obtain one for its floored counterpart $\lfloor\delta^{\alpha}_{\min}\rfloor(N)$ (both defined in Section 1) for real sequences $(a_{n})_{n\geq 1}$ of distinct numbers. Our theorems recover [ABM, Theorems 1–3], as well as the result from [ABM, Section 4.3]. Furthermore, we establish lower bounds for the minimal gaps of well-spaced sequences and, more generally, of a broader family that contains them.
Key words and phrases:
Metric Diophantine approximation, Duffin–Schaeffer conjecture, minimal gap of sequences mod $1$, difference set.
2020 Mathematics Subject Classification:
Primary 11B05, 11J71, 11J83, 11M06; Secondary 11K06, 11K38
1. Introduction
The theory of uniform distribution modulo one is a fundamental area at the intersection of number theory, ergodic theory, and metric Diophantine approximation. A central problem within this field is to understand the local statistics of the sequences $(\alpha a_{n})_{n\geq 1}$, where $\alpha$ is a real parameter and $(a_{n})_{n\geq 1}$ is a given sequence of distinct real numbers. A key statistic capturing the finest scale behaviour is the minimal gap, defined for $N$ terms as
(1) $\delta_{\min}^{\alpha}(N)=\min\{\|\alpha(a_{m}-a_{n})\|:1\leq m\neq n\leq N\},$
where $\|x\|=\displaystyle\min_{k\in\mathbb{Z}}|x-k|$. The behaviour of $\delta_{\min}^{\alpha}(N)$ as $N\to\infty$ for generic $\alpha$ (that is, for almost all $\alpha$ with respect to the Lebesgue measure) reveals the deep arithmetic properties of the sequence $(a_{n})_{n\geq 1}$. As direct computation of the minimal gap $\delta_{\min}^{\alpha}(N)$ for fixed $\alpha$ is typically difficult, we analyse its metric properties, focusing on the almost-sure behaviour and distribution.
The expected behaviour of the minimal gap serves as a powerful tool for distinguishing random from deterministic sequences. The minimal gap for a random sequence is always at most $1/N$, as $1/N$ is the average spacing. For a sequence of $N$ independent and uniformly distributed random points in the unit interval, the minimal gap is almost surely of size about $1/N^{2}$, a hallmark of Poissonian statistics where points exhibit no correlations. This random benchmark provides a crucial baseline: a deterministic sequence is considered to exhibit Poissonian behaviour if its minimal gap also decays as $1/N^{2}$ as in the random case.
In contrast to the random case, more structured sequences, such as $(\sqrt{n})_{n\notin\square}$, are more rigidly spaced, resulting in a larger minimal gap that decays more slowly (see [Regavim]), thus signifying a clear departure from random behaviour. The minimal gap in this case is known to be of the order $2/N^{3/2}$, which is much larger than the random scaling of $1/N^{2}$. Since polynomial sequences $(n^{d})_{n\geq 1}$ for $d\geq 2$ and lacunary sequences have nearly minimal additive energy (close to $N^{2}$), the minimal gap $\delta_{\min}^{\alpha}(N)$ is of order $1/N^{2}$ up to sub-power factors. This follows from the following results of Rudnick [Rudnick, Theorems 1–2].
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct integers, and let its additive energy be nearly minimal, that is, $\ll N^{2+o(1)}$. Then, for almost all $\alpha$, any $\eta>0$ , and for all sufficiently large $N$, we have
(2) $\frac{1}{N^{2+\eta}}<\delta_{\min}^{\alpha}(N)<\frac{1}{N^{2-\eta}}.$
The landmark work of Aistleitner, El-Baz, and Munsch [ABM] established that the metric theory of minimal gaps for integer sequences is not controlled by the additive energy, but rather by the cardinality of the difference set $\{a_{i}-a_{j}:1\leq i\neq j\leq N\}$, providing a comprehensive framework that connects the almost-sure minimal gap of $(\alpha a_{n})_{n\geq 1}$ directly to its difference set structure. This article significantly extends their line of research to sequences $(a_{n})_{n\geq 1}$ of arbitrary distinct positive real numbers. We introduce and study several natural variants of the minimal gap $\delta_{\min}^{\alpha}(N)$, namely $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$, $\widetilde{\delta}_{\min}^{\alpha}(N)$, and $\widehat{\delta}_{\min}^{\alpha}(N)$ (defined in Sections 1 and 3). A primary aim is to understand the relationships between these quantities and to determine their typical asymptotic order for almost all $\alpha$.
Our main results are multifaceted. First, we establish a general lower bound and a weak upper bound for $\delta_{\min}^{\alpha}(N)$ under the mild assumption that the sequence $(a_{n})_{n\geq 1}$ has a uniform positive lower bound on its gaps, $\inf_{m\neq n}|a_{m}-a_{n}|\geq c>0$. These bounds are expressed in terms of $D_{N}$, defined as the number of distinct positive differences between $\{a_{1},\ldots,a_{N}\}$, and feature precise logarithmic factors endemic to such metric problems.
Second, we prove a more refined theorem where the constant lower bound $c$ is replaced by a monotonically decreasing function $c(N)$, allowing the treatment of sequences whose terms can cluster. This provides a framework for handling a wider class of sequences, including polynomial sequences with exponents less than $1$.
Third, we establish several upper bounds for $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$, where
(3) $\lfloor\delta_{\min}^{\alpha}\rfloor(N)=\min\left\{\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|:1\leq m\neq n\leq N\right\},$
using the powerful machinery of the Duffin-Schaeffer conjecture, recently resolved by Koukoulopoulos and Maynard (see [KoukouMaynard]). These include three types of bounds: one that holds for infinitely many $N$ and depends on $H_{N}$ (the cardinality of the set of floored positive differences) and the size of $a_{N}$, another that depends solely on $H_{N}$, and a third bound in terms of $H_{N}$ that is valid for all sufficiently large $N$.
Finally, for any given sequence $(\eta(N))_{N\geq 1}$ of non-negative reals, we present a complete zero-one law that characterises whether the inequality $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$ holds for infinitely many $N$ and for almost no or almost all $\alpha$. This theorem provides a unifying criterion that recovers and generalises previous results.
Our work demonstrates that the fundamental principles governing the fine-scale statistics of sequences, previously understood primarily in the integer setting, have powerful extensions to the realm of real sequences. The results have applications to lacunary sequences, Beatty sequences, and sequences defined by smooth, sufficiently monotonic functions, several examples of which are explored in detail in later sections.
The investigation of sequences modulo one sits at the rich confluence of number theory, ergodic theory, and probability. A sequence $(a_{n})_{n\geq 1}$ in $[0,1)$ is considered uniformly distributed if, in the limit, the proportion of its terms that fall into any subinterval is equal to the length of the interval. This “global” equidistribution, a concept firmly established since the pioneering work of Weyl [Weyl1916], is elegantly quantified by discrepancy theory, which measures the maximum deviation from this ideal distribution [MichaelTichy, kuipers1974uniform].
However, global uniformity often masks a more complex microscopic reality. Two sequences can be perfectly uniformly distributed, yet exhibit radically different local behaviours. The true arithmetic character of a sequence is frequently revealed not by its global distribution but by its local statistics, the fine-scale patterns such as the distribution of gaps between consecutive points and the correlation between the spacings of multiple points. The study of these statistics, including pair correlation and nearest-neighbour gap distributions, has proven to be a remarkably sensitive probe. It forges deep and often unexpected connections to diverse fields such as random matrix theory, quantum chaos, and the metric Diophantine approximation, providing a unifying language for describing randomness and structure in deterministic systems.
Although local statistics of integer sequences have been extensively studied (see [ABM2021, Aistleitner-Larcher-Lewko, Bloom-Chow, Rudnick, Rudnick-Sarnak]), with important work on the distribution of spacings and pair correlation for sequences like $(n^{2}\alpha)_{n\geq 1}$ [RudnickSarnakAlexandru2001] and other lacunary sequences [RudnickAlexandru1999, RudnickAlexandru2002], the natural and vast generalisation to sequences of arbitrary real numbers has remained largely unexplored. This gap is significant, as many sequences of central importance in analysis and mathematical physics—such as lacunary sequences, values of smooth functions, and geometric progressions with non-integer ratios—reside firmly within this real-valued realm. A comprehensive theory for real sequences is, therefore, not merely a technical extension but a necessary step for a complete understanding.
This article is motivated by the goal of establishing such a theory for the minimal gap statistic. We seek to generalise the seminal findings of Aistleitner, El-Baz, and Munsch [ABM] for integer sequences, which tied the minimal gap to the cardinality of the difference set, into the domain of real sequences. By introducing and analysing new, natural variants of the minimal gap tailored for real numbers, we aim to uncover the fundamental principles that govern their fine-scale distribution. Our work provides a unified framework that not only recovers classical results but also opens the door to analysing a much broader and richer class of sequences, thereby deepening our understanding of the interplay between arithmetic structure and random-like behaviour.
1.1. Organisation of the article
The article is structured as follows. In Section 2, we summarise all main results of this article without proofs. We also highlight key corollaries and provide a preview of applications to various sequence families.
Section 3 is dedicated to the study of minimal gap statistics for sequences of distinct positive real numbers. We begin by defining four key metrics: $\delta_{\min}^{\alpha}(N)$, $\widetilde{\delta}_{\min}^{\alpha}(N)$, $\widehat{\delta}_{\min}^{\alpha}(N)$, and $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$. We then establish the fundamental chain of inequalities $\delta_{\min}^{\alpha}(N)\leq\widetilde{\delta}_{\min}^{\alpha}(N)\leq\widehat{\delta}_{\min}^{\alpha}(N)$. The core of this section is the proof of Theorem 1, which provides nearly optimal lower and upper bounds for $\delta_{\min}^{\alpha}(N)$ for sequences with uniform positive separation $\inf_{m\neq n}|a_{m}-a_{n}|\geq c>0$. The proof employs a careful application of the Borel–Cantelli lemma. We also survey relevant prior results, particularly those of Aistleitner, El-Baz, and Munsch for integer sequences (see [ABM]).
We then generalise this result in Theorem 5 to sequences where the minimal separation $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|$ is bounded below by a monotonically decreasing sequence. A novel and careful enumeration of the difference set is introduced to handle this more delicate situation. A key consequence, detailed in Corollary 23, recovers and extends known results, such as Rudnick’s theorem, by showing that for sequences with a fixed positive separation, one has $\delta_{\min}^{\alpha}(N)\gg 1/N^{2+\eta}$ for any $\eta>0$ and for almost all $\alpha$.
In Section 4, we establish several key upper bounds for the metric $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$. First, we prove a general upper bound in Theorem 6. Subsequently, we refine this analysis to prove Theorem 9, which provides an upper bound for $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$ that holds for infinitely many $N$ and, crucially, depends only on the counting function $H_{N}$. A central component of this proof is a direct application of the celebrated Koukoulopoulos–Maynard theorem (Theorem 35). Finally, in Section 4.4, we establish Theorem 11, which gives an upper bound for $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$ in terms of $H_{N}$ that is valid for all sufficiently large $N$.
In Section 5, we present a complete zero-one law characterisation for $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$ to occur infinitely often in Theorem 13. The proofs of Lemma 49, Theorem 9, and Theorem 13 build upon the corresponding arguments in [ABM] for integer sequences, incorporating key modifications developed in previous lemmas to adapt to our setting. We provide detailed proofs for completeness
2. Main Results
Before stating the main results, we collect the key notation and definitions used throughout the article.
2.1. Notation and definitions
  • $\lfloor x\rfloor$: The floor of a real number $x$, denoting the greatest integer less than or equal to $x$.
  • $\{x\}$: The fractional part of a real number $x$, defined by $\{x\}=x-\lfloor x\rfloor$.
  • $\|x\|$: The distance from $x$ to the nearest integer. For $x\in\mathbb{R}$, $\|x\|=\min(\{x\},1-\{x\})$.
  • $\log x$: $\max(1,\log x)$, where $\log$ is the natural logarithm.
  • $\log_{k}x$: The $k$-th iterated logarithm, defined recursively by $\log_{1}x=\log x$ and $\log_{k}x=\log(\log_{k-1}x)$ for $k\geq 2$. Thus, $1/\log n$, $\log_{2}n$, $\log_{3}n$, etc. are positive and well-defined for all integers $n\geq 1$.
  • $f(x)\ll g(x)$: $|f(x)|\leq C|g(x)|$ for some constant $C>0$ and for all $x$ in a specified domain. Equivalently, $f(x)=O(g(x))$. We write $g(x)\gg f(x)$ to mean $f(x)\ll g(x)$.
  • $f(x)=o(g(x))$: $\frac{f(x)}{g(x)}\to 0$ as $x\to\infty$.
  • $\nu_{p}(n)$: The $p$-adic valuation of $n\in\mathbb{N}$, that is, the largest integer $k$ such that $p^{k}$ divides $n$.
  • $A^{+}$: The set of all positive elements of a set $A$.
  • $\lfloor A\rfloor$: For any subset $A\subseteq\mathbb{R}$, $\lfloor A\rfloor=\{\lfloor a\rfloor:a\in A\}$.
  • $A_{N}$: The finite set of the first $N$ terms, $A_{N}=\{a_{1},a_{2},\dots,a_{N}\}$.
  • $A_{N}-A_{N}$: The entire difference set, $\{a_{m}-a_{n}:1\leq m,n\leq N\}$.
  • $(A_{N}-A_{N})^{+}$: The set of positive differences, $\{|a_{m}-a_{n}|:1\leq m\neq n\leq N\}$.
  • $\lfloor(A_{N}-A_{N})^{+}\rfloor$: The set of floored positive differences.
  • $(\lfloor A_{N}-A_{N}\rfloor)^{+}$: The set of positive floored differences.
  • $C_{N}$,  $D_{N}$: $C_{N}=\#(A_{N}-A_{N})$,  $D_{N}=\#(A_{N}-A_{N})^{+}$.
  • $H_{N}$,  $T_{N}$: $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$,  $T_{N}=\#(\lfloor A_{N}-A_{N}\rfloor)^{+}$.
  • $\mathcal{N}(k)$: The first index $M$ at which the positive integer $k$ appears in the set of floored positive differences, defined by
    (4) $\mathcal{N}(k)=\min\left\{M\geq 1:k\in\lfloor(A_{M}-A_{M})^{+}\rfloor\right\}.$
    If $k\notin\cup_{N\geq 1}\lfloor(A_{N}-A_{N})^{+}\rfloor$, we set $\mathcal{N}(k)=\infty$ and adopt the convention that the supremum of the empty set is zero.
  • Almost all $\alpha$: A property that holds for all $\alpha$ in a subset of $[0,1]$ with the Lebesgue measure $1$.
  • $\mathds{1}_{E}(*)$: The indicator function of $E$, that is, $\mathds{1}_{E}(x)$ is equal to 1 if $x\in E$ and 0 otherwise. We adopt the convention that $\mathds{1}(x=0)$ denotes $\mathds{1}_{\{0\}}(x)$.
Equipped with this notation, we now state our main results. This work establishes new results on the fine-scale statistics of sequences $(\alpha a_{n})_{n\geq 1}$ for generic $\alpha$. Our theorems provide comprehensive bounds, both lower and upper, that hold for infinitely many $N$ and for all sufficiently large $N$, along with a precise zero-one law for the minimal gap and its variants. The results reveal how the distribution of gaps depends on the arithmetic structure of the sequence through the cardinalities $D_{N}$, $H_{N}$, and related quantities.
2.2. General lower and upper bounds
Our first main result provides nearly optimal lower and upper bounds for the minimal gap $\delta_{\min}^{\alpha}(N)$ for sequences with a positive uniform separation. The limits are given in terms of the quantity $D_{N}$, the number of distinct positive differences between the first $N$ terms.
Theorem 1.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that $\inf_{m\neq n}|a_{m}-a_{n}|\geq c$ for some positive constant $c$. Let $\epsilon>0$, and let $D_{N}$ be defined as in Section 2.1. Then, for almost all $\alpha\in[0,1]$, and for all sufficiently large NN ((that is, there exists $N_{0}(\alpha)$ such that for all $N\geq N_{0}(\alpha)$)), we have
$\delta_{\mathrm{min}}^{\alpha}(N)\geq\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}(N)\leq 1-\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}.$
Remark 2.
Since $D_{N}\leq C_{N}$, the statement of Theorem 1 also holds when $D_{N}$ is replaced by $C_{N}$.
Remark 3.
In particular, condition $\inf_{m\neq n}|a_{m}-a_{n}|\geq 1$ is automatically satisfied for any sequence $(a_{n})_{n\geq 1}$ of distinct positive integers. Therefore, we recover the second part of [ABM, Theorem 1].
Remark 4.
For any $\epsilon>0$, we have $\frac{\log_{2}N}{\log{N}}+(1+\epsilon)\frac{\log_{3}N}{\log{N}}=o(1)$ as $N\to\infty$. Corresponding to a fixed $\eta>0$, there exists $N_{0}^{{}^{\prime}}(\eta)$ such that $\frac{\log_{2}N+\log_{3}N^{(1+\epsilon)}}{\log{N}}<\eta$, that is, $\log N\log_{2}N^{(1+\epsilon)}<N^{\eta}$, for all $N\geq N_{0}^{{}^{\prime}}(\eta)$. Let $N_{0}^{{}^{\prime}}(\alpha)$ be such that Theorem 1 holds for all $N$ larger than $N_{0}^{{}^{\prime}}(\alpha)$. Define $N_{0}(\alpha):=\max(N_{0}^{{}^{\prime}}(\alpha),N_{0}^{{}^{\prime}}(\eta))$. Since $D_{N}\leq C_{N}\leq N^{2}$, we have
(5) $\delta_{\mathrm{min}}^{\alpha}(N)\geq\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\geq\frac{1}{N^{2}}\frac{1}{N^{\eta}}=\frac{1}{N^{2+\eta}}$
for almost all $\alpha\in[0,1]$ and for every $N\geq N_{0}(\alpha)$.
In other words, for each $\eta>0$ and for almost all $\alpha\in[0,1]$, we have $\delta_{\mathrm{min}}^{\alpha}(N)\geq\frac{1}{N^{2+\eta}}$ for every $N\geq N_{0}(\alpha)$. Thus, we recover [Rudnick, Theorem 1] for a sequence of distinct real numbers.
2.3. Refined bounds for sequences with decaying gaps
We next present a refinement of Theorem 1 that applies to sequences where the minimal gap between distinct terms is allowed to decay to zero at a controlled rate.
Theorem 5.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq c(N)$, where $(c(N))_{N\geq 1}$ is a monotone decreasing sequence of positive reals. Let $\epsilon>0$, and let $D_{N}$ be defined as in Section 2.1. Then, for almost all $\alpha\in[0,1]$, and for all sufficiently large $N$, we have
(a) $\delta^{\alpha}_{\min}(N)\geq\dfrac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$, and
(b) $\delta^{\alpha}_{\min}(N)\leq 1-\dfrac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$.
2.4. Upper bound via the Duffin-Schaeffer conjecture
For the quantity $\lfloor\delta^{\alpha}_{\min}\rfloor(N)$, we prove three types of upper bounds: first, a bound that holds for infinitely many $N$ and depends on the size of $a_{N}$; second, a bound that is independent of the size of $a_{N}$ and instead relies only on $H_{N}$; third, an upper bound for $\lfloor\delta_{\min}^{\alpha}\rfloor(N)$ in terms of $H_{N}$ that is valid for all sufficiently large $N$.
Theorem 6.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Then, for almost all $\alpha\in[0,1]$, we have
(6) $\lfloor\delta^{\alpha}_{\min}\rfloor(N)\leq\frac{4\log_{2}B_{N}}{H_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}\quad\text{for infinitely many }N,$
where $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\displaystyle\min_{1\leq m\neq n\leq N}\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|$, $B_{N}=\displaystyle\max_{1\leq m\neq n\leq N}\lfloor|a_{m}-a_{n}|\rfloor$ and $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$.
Remark 7.
Since $B_{N}\leq\lfloor\max(a_{1},\ldots,a_{N})\rfloor$, the statement of Theorem 6 also holds when $B_{N}$ is replaced by $\lfloor\max(a_{1},\ldots,a_{N})\rfloor$.
Remark 8.
For a sequence of distinct positive integers, we have $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq 1$, $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\delta_{\min}^{\alpha}(N)$, and $D_{N}=H_{N}$. As a consequence, Theorem 6 yields a result of Aistleitner, El-Baz, and Munsch [ABM, Theorem 1 (Equation 5)], which corresponds to the integer case.
Theorem 9.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Then, for almost all $\alpha\in[0,1]$, we have
(7) $\lfloor\delta^{\alpha}_{\min}\rfloor(N)\leq\frac{1}{H_{N}}\quad\text{for infinitely many }N.$
Remark 10.
Theorem 9 finds a natural counterpart in the following result, which establishes a similar upper bound for $\lfloor\delta^{\alpha}_{\min}\rfloor(N)$, valid for all sufficiently large $N$.
Theorem 11.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Then, for every $\epsilon>0$ and almost all $\alpha$,
$\lfloor\delta^{\alpha}_{\min}\rfloor(N)\leq\frac{N^{\epsilon}}{H_{N}}\quad\text{for all except finitely many }N.$
Remark 12.
When $(a_{n})_{n\geq 1}$ is an integer sequence, we have $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\delta_{\min}^{\alpha}(N)$, and $D_{N}=H_{N}$. Thus, Theorems 9 and 11 recover the results in [ABM, Section 4.3] and [ABM, Theorem 2], respectively.
2.5. A zero-one law for the minimal gap
Our final main result is a comprehensive zero-one law that completely characterises the set of $\alpha$ for which $\lfloor\delta^{\alpha}_{\min}\rfloor(N)$ is infinitely often smaller than a given threshold function $\eta(N)$.
Theorem 13.
Let $\left(a_{n}\right)_{n\geq 1}$ be a sequence of positive real numbers such that $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq 1$. Let $(\eta(N))_{N\geq 1}$ be a sequence of non-negative reals. Let $\mathcal{A}$ denote the set of those $\alpha\in[0,1]$ for which $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$ holds for infinitely many $N$. Then, we have $\lambda(\mathcal{A})=0$ or $\lambda(\mathcal{A})=1$, depending on whether the series
(8) $\sum_{k=1}^{\infty}\varphi(k)\sup_{b\geq 1}\left\{\frac{\sup_{\ell\geq\mathcal{N}(bk)}\eta(\ell)}{bk}\right\}$
is convergent or divergent, respectively, where $\varphi$ is Euler’s totient function and $\lambda$ denotes the Lebesgue measure on $\mathbb{R}$.
Remark 14.
Under the assumption that $\eta$ is decreasing, the series for which convergence or divergence is to be checked in Theorem 13 takes the following simpler form
$\sum_{k=1}^{\infty}\varphi(k)\sup_{b\geq 1}\left\{\frac{\eta(\mathcal{N}(bk))}{bk}\right\}.$
Remark 15.
Theorem 13 provides a sharp criterion reducing the problem of determining the almost sure behaviour of the minimal gap of $(\alpha a_{n})_{n\geq 1}$ to the arithmetic structure of $\lfloor(A_{N}-A_{N})^{+}\rfloor$, showing that this floored difference set alone determines the typical asymptotic order, at least in principle.
Remark 16.
When $(a_{n})_{n\geq 1}$ is a sequence of distinct positive integers, we have $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq 1$ and $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\delta_{\min}^{\alpha}(N)$. Thus, we recover [ABM, Theorem $3$] from Theorem 13.
2.6. Applications and examples
To illustrate the applicability of our theorems, we present several classes of sequences for which our methods yield new estimates for $\delta_{\mathrm{min}}^{\alpha}(N)$.
  • Polynomial sequences: For $a_{n}=n^{\theta}$ with $\theta\geq 1$, we get $\delta_{\mathrm{min}}^{\alpha}(N)\gg(N^{2}\log N(\log_{2}N)^{1+\epsilon})^{-1}$ for any $\epsilon>0$.
  • Polynomial sequences with exponent less than one: For $a_{n}=n^{\theta}$ with $\theta<1$, we obtain $\delta_{\mathrm{min}}^{\alpha}(N)\gg(N^{3-\theta}\log N(\log_{2}N)^{1+\epsilon})^{-1}$ for any $\epsilon>0$.
  • Lacunary sequences: Sequences that satisfy $\frac{a_{n+1}}{a_{n}}\geq D>1$, such as $a_{n}=e^{n}$ or $\pi^{n}$.
  • Beatty sequences: Sequences of the form $a_{n}=\lfloor n\theta\rfloor$ for irrational $\theta>1$.
  • Sequences in arithmetic progressions: Sequences of the form $a_{n}=a+(n-1)d$ where $a$ and $d$ are positive constants.
  • Sequences under smooth maps: If $f$ is differentiable with $|f^{\prime}|$ bounded away from zero, then $(f(a_{n}))_{n\geq 1}$ inherits the minimal gap properties of $(a_{n})_{n\geq 1}$.
The subsequent sections of this article are devoted to the detailed proofs of the results mentioned in this section, followed by a discussion of their implications and connections to the existing literature.
3. minimal gap statistics
For a given $\alpha\in[0,1]$ and a sequence $(a_{n})_{n\geq 1}$ of distinct real numbers, let $\delta_{\min}^{\alpha}(N)$ and $\lfloor\delta^{\alpha}_{\min}\rfloor(N)$ be defined as in Section 1. We now introduce the following related variants of the minimal gap:
  • $\widetilde{\delta}_{\min}^{\alpha}(N)$: The minimal fractional part of scaled differences,
    (9) $\widetilde{\delta}_{\min}^{\alpha}(N)=\min\left\{\{\alpha|a_{m}-a_{n}|\}:1\leq m\neq n\leq N\right\}.$
  • $\widehat{\delta}_{\min}^{\alpha}(N)$: The minimal scaled absolute difference,
    (10) $\widehat{\delta}_{\min}^{\alpha}(N)=\min\left\{\alpha|a_{m}-a_{n}|:1\leq m\neq n\leq N\right\}.$
Remark 17.
The expression $\delta_{\min}^{\alpha}(N)$ depends on the sequence $(a_{n})_{n\geq 1}$, as do the quantities $\widetilde{\delta}_{\min}^{\alpha}(N)$, $\widehat{\delta}_{\min}^{\alpha}(N)$, and $\lfloor\delta^{\alpha}_{\min}\rfloor(N)$, but we suppress this dependence in the notation for simplicity.
Note that
(11) $\widetilde{\delta}_{\min}^{\alpha}(N)=$ $\min\left\{\{\alpha|a_{m}-a_{n}|\}:1\leq m\neq n\leq N\right\}$
(12) $\leq$ $\{\alpha|a_{m^{\prime}}-a_{n^{\prime}}|\}$
(13) $\leq$ $\alpha|a_{m^{\prime}}-a_{n^{\prime}}|\quad\text{for all }m^{\prime}\neq n^{\prime}\leq N.$
Now, taking the minimum over all $m^{\prime}\neq n^{\prime}\leq N$, we obtain $\widetilde{\delta}_{\min}^{\alpha}(N)\leq\widehat{\delta}_{\min}^{\alpha}(N)$.
Similarly,
(14) $\delta_{\min}^{\alpha}(N)=$ $\min\left\{\|\alpha(a_{m}-a_{n})\|:1\leq m\neq n\leq N\right\}$
(15) $\leq$ $\|\alpha(a_{m^{\prime}}-a_{n^{\prime}})\|$
(16) $=$ $\|\alpha|a_{m^{\prime}}-a_{n^{\prime}}|\|$
(17) $\leq$ $\{\alpha|a_{m^{\prime}}-a_{n^{\prime}}|\}\quad\text{for all }m^{\prime}\neq n^{\prime}\leq N.$
Now, taking the minimum over all $m^{\prime}\neq n^{\prime}\leq N$, we obtain $\delta_{\min}^{\alpha}(N)\leq\widetilde{\delta}_{\min}^{\alpha}(N)$. Therefore, for all $N\in\mathbb{N}$, we obtain
(18) $\delta_{\min}^{\alpha}(N)\leq\widetilde{\delta}_{\min}^{\alpha}(N)\leq\widehat{\delta}_{\min}^{\alpha}(N).$
Remark 18.
It is evident from (18) that when studying the lower bounds for such quantities, it suffices to find a lower bound for $\delta_{\min}^{\alpha}(N)$ (which will provide a lower bound for the other two, possibly a weaker one), and to find an upper bound it suffices to find an upper bound for the quantity $\widehat{\delta}_{\min}^{\alpha}(N)$. Determining the precise growth rates of these three quantities poses interesting questions for further study.
Remark 19.
If $(a_{n})_{n\geq 1}$ is a sequence of positive integers, then we have $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\delta_{\min}^{\alpha}(N)$.
3.1. Proof of Theorem 1
We first arrange the elements of $A=\displaystyle\bigcup_{N\geq 1}\left(A_{N}-A_{N}\right)^{+}$ in a sequence. Let $\left(z_{n}\right)_{n\geq 1}$ be a sequence of distinct positive real numbers such that for all $N$,
(19) $\left\{z_{n}:1\leq n\leq D_{N}\right\}=\left(A_{N}-A_{N}\right)^{+},$
with $(0=)D_{1}<D_{2}<\ldots<D_{N}<D_{N+1}<\ldots$, and $D_{N}\geq(N-1)$.
The sequence $\left(z_{n}\right)_{n\geq 1}$ need not be unique, as $D_{N+1}-D_{N}$ can be as large as $N$. However, the same proof works for any choice of $\left(z_{n}\right)_{n\geq 1}$ that satisfies (19). Since $(a_{n})_{n\geq 1}$ is a sequence of distinct positive reals satisfying $\inf_{m\neq n}|a_{m}-a_{n}|\geq c$, we have $z_{n}\geq c$ for all $n$, and hence $\lfloor z_{n}\rfloor\geq 0$ for all $n$.
For $n\geq 1$, we set
(20) $\psi(n)=\frac{1}{n(\log\sqrt{n})\left(\log_{2}\sqrt{n}\right)^{1+\varepsilon}},$
$J_{n}(a)=\left(\frac{a}{z_{n}}-\frac{\psi(n)}{z_{n}},\frac{a}{z_{n}}+\frac{\psi(n)}{z_{n}}\right),$
and
(21) $S_{n}=[0,1]\cap\Big(\bigcup_{0\leq a\leq\lfloor z_{n}\rfloor+1}J_{n}(a)\Big).$
We have
(22) $\sum_{n=1}^{\infty}\lambda\left(S_{n}\right)\leq$ $\sum_{n=1}^{\infty}\min\left\{\frac{2\psi(n)}{z_{n}}(\lfloor z_{n}\rfloor+2),1\right\}$
(23) $\leq$ $\sum_{n=1}^{\infty}2\psi(n)\left(1+\frac{2}{z_{n}}\right)$
(24) $\leq$ $\sum_{n=1}^{\infty}2\left(1+\frac{2}{c}\right)\psi(n)\leq\sum_{n=1}^{\infty}\frac{2\left(1+\frac{2}{c}\right)}{n(\log\sqrt{n})\left(\log_{2}\sqrt{n}\right)^{1+\varepsilon}}<\infty.$
Note that since $S_{n}\subseteq[0,1]$ we have $\lambda(S_{n})=\text{Prob}(\{\alpha\in[0,1]:\alpha\in S_{n}\})$, and hence
$\sum_{n\geq 1}\text{Prob}(\{\alpha\in[0,1]:\alpha\in S_{n}\})<\infty.$
By the first Borel-Cantelli lemma, since $\sum_{n=1}^{\infty}\lambda(S_{n})<\infty$, we have
(25) $0=$ $\text{Prob}(\{S_{n}\text{ infinitely often}\})=\text{Prob}\Big(\{\alpha\in[0,1]:\alpha\in\bigcap_{n=1}^{\infty}\bigcup_{i=n}^{\infty}S_{i}\}\Big).$
In other words, $\text{Prob}\Big(\{\alpha\in[0,1]:\alpha\notin\displaystyle\bigcap_{n=1}^{\infty}\displaystyle\bigcup_{i=n}^{\infty}S_{i}\}\Big)=1$, that is, there exists a set $\tilde{Q}\subseteq[0,1]$ with $\lambda(\tilde{Q})=1$ such that for all $\alpha\in\tilde{Q}$, we have $\alpha\notin\displaystyle\bigcap_{n=1}^{\infty}\bigcup_{i=n}^{\infty}S_{i}$. This means that there exists $N_{1}(\alpha)\in\mathbb{N}$ such that $\alpha\notin\displaystyle\bigcup_{i=N_{1}(\alpha)}^{\infty}S_{i}$, that is, $\alpha\notin S_{n}$ for all $n\geq N_{1}(\alpha)$.
Therefore, for all $n\geq N_{1}(\alpha)$, we have
(26) $\alpha\notin[0,1]\cap\Big(\bigcup_{0\leq a\leq\lfloor z_{n}\rfloor+1}J_{n}(a)\Big)$
(27) $\overset{\alpha\in[0,1]}{\implies}$ $\alpha\notin\bigcup_{0\leq a\leq\lfloor z_{n}\rfloor+1}J_{n}(a)$
(28) $\implies$ $|\alpha-\frac{a}{z_{n}}|\geq\frac{\psi(n)}{z_{n}}\text{ for all }a\in\{0,1,\ldots,\lfloor z_{n}\rfloor+1\}$
(29) $\implies$ $|\alpha z_{n}-a|\geq\psi(n)\text{ for all }a\in\{0,1,\ldots,\lfloor z_{n}\rfloor+1\}.$
Therefore, for all $\alpha\in\tilde{Q}$ and all $n\geq N_{1}(\alpha)$, we have
(30) $|\alpha z_{n}-a|\geq\psi(n)\text{ for all }a\in\{0,1,\ldots,\lfloor z_{n}\rfloor+1\}.$
Case 1: Let $z_{n}\in[c,1)$. Then $\lfloor z_{n}\rfloor=0$. We have $0\leq\alpha z_{n}\leq z_{n}=\lfloor z_{n}\rfloor+\{z_{n}\}=\{z_{n}\}<1$, so $|\alpha z_{n}-1|=1-\alpha z_{n}=1-\{\alpha z_{n}\}$. Therefore, using (30), we obtain
(31) $\|\alpha z_{n}\|=\min(\{\alpha z_{n}\},1-\{\alpha z_{n}\})=$ $\displaystyle\min_{a\in\{0,1\}}|\alpha z_{n}-a|=\displaystyle\min_{a\in\{0,\ldots,\lfloor z_{n}\rfloor+1\}}|\alpha z_{n}-a|\geq\psi(n).$
Case 2: Let $z_{n}\in[1,\infty)$. Then, for all $\alpha\in[0,1]$, we have $0\leq\alpha z_{n}\leq z_{n}<\lfloor z_{n}\rfloor+1$. Therefore, using (30), we obtain
[…]
Therefore, combining (31) and (32), we find that for all $\alpha\in\tilde{Q}$ and all $n\geq N_{1}(\alpha)$, we have $\alpha z_{n}||\geq\psi(n)$.
(a) Let $a_{m},a_{n}\in A_{N}$ with $m\neq n$. Then, $|a_{m}-a_{n}|=z_{n^{\prime}}$ for some $n^{\prime}=n^{\prime}(m,n)\leq D_{N}$, where $D_{N}\leq N^{2}$.
We now prove that for all sufficiently large $N$ and for any $\epsilon>0$, $\|\alpha z_{n^{\prime}}\|>\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$. We consider a finite set of indices $\{1,2,\dots,N_{1}(\alpha)-1\}$. Since $z_{n}$ are positive and distinct, and since $\alpha$ is such that $\alpha z_{n}\notin\mathbb{Z}$ for any $n$ (which is true for almost all $\alpha$), there exists a constant $\eta(\alpha)>0$ such that
(33) $\|\alpha z_{n}\|\geq\eta(\alpha)\quad\text{for all}\quad n<N_{1}(\alpha).$
Furthermore, since $D_{N}\geq N-1$, we have
(34) $\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\to 0\quad\text{as}\quad N\to\infty.$
Therefore, there exists $N_{2}(\alpha)$ such that for all $N\geq N_{2}(\alpha)$,
(35) $\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}<\eta(\alpha).$
Define $N_{0}(\alpha)=\max(N_{1}(\alpha),N_{2}(\alpha))$. Then, for all $N\geq N_{0}(\alpha)$ and every pair of distinct elements $a_{m},a_{n}\in A_{N}$ with corresponding index $n^{\prime}$:
  • If $n^{\prime}\geq N_{1}(\alpha)$, then using the facts that $\psi$ is decreasing and $n^{\prime}\leq D_{N}$, we have
    (36) $\|\alpha z_{n^{\prime}}\|\geq\psi(n^{\prime})\geq\psi(D_{N})=\frac{1}{D_{N}(\log\sqrt{D_{N}})(\log_{2}\sqrt{D_{N}})^{1+\epsilon}}\geq\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}},$
    where we used the facts that $D_{N}\leq C_{N}\leq N^{2}$.
  • If $n^{\prime}<N_{1}(\alpha)$, then
    (37) $\|\alpha z_{n^{\prime}}\|\geq\eta(\alpha)>\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}.$
Therefore, for all $N\geq N_{0}(\alpha)$, and all distinct $a_{m},a_{n}\in A_{N}$, we have
[…]
for almost all $\alpha\in[0,1]$. The proof now follows by taking the minimum over all $a_{m},a_{n}\in A_{N}$.
(b) Since $0\leq\alpha z_{n}\leq z_{n}$, we have $0\leq\lfloor\alpha z_{n}\rfloor\leq\lfloor z_{n}\rfloor.$ In particular, taking $a=\lfloor\alpha z_{n}\rfloor$ in (30), we have the following for all $\alpha\in\tilde{Q}$ (with $\lambda(\tilde{Q})=1$) and all $n\geq N_{0}(\alpha)$,
(39) $\{\alpha z_{n}\}\geq\psi(n).$
Therefore, for all $N\geq N_{0}(\alpha)$ and all $a_{m},a_{n}\in A_{N}$ with ${m}\neq{n}$, the inequality
(40) $\left\{\alpha\left|a_{m}-a_{n}\right|\right\}\geq\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$
holds for almost all $\alpha\in[0,1]$.
Note that t= $t||=\text{ min }(\{t\},1-\{t\})\leq 1-\{$t\} and t=t||-t||=||t|| for any $t\in\mathbb{R}$. Therefore, using (39), for all $N\geq N_{0}(\alpha)$ and all distinct $a_{m},a_{n}\in A_{N}$, we have
(41) $\left\|\alpha\left(a_{m}-a_{n}\right)\right\|=\left\|\alpha|a_{m}-a_{n}|\right\|\leq 1-\{\alpha|a_{m}-a_{n}|\}\leq 1-\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$
for almost all $\alpha\in[0,1]$. The proof now follows by taking the minimum over all $a_{m},a_{n}\in A_{N}$.
Example 20.
Let $\epsilon>0$, $m\in\mathbb{N}$, and $d\geq 0$ be a fixed constant. Then, Theorem 1 implies that for any integer sequence $(a_{n})_{n\geq 1}$, the sequence $(\frac{a_{n}}{m}+d)_{n\geq 1}$ satisfy
$\delta_{\mathrm{min}}^{\alpha}(N)\geq\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$
for almost all $\alpha\in[0,1]$, and for all sufficiently large $N$, where $D_{N}$ is the cardinality of the set of positive differences for the sequence $(\frac{a_{n}}{m}+d)_{n\geq 1}$.
Corollary 21.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that $\inf_{m\neq n}|a_{m}-a_{n}|\geq c$ for some constant $c>0$. Let $\epsilon>0$, and let $f$ be a differentiable function on $\inf_{n}a_{n},\infty)$, which is positive-valued on $(a_{n})_{n\geq 1}$ and for which there exists $\delta_{0}>0$ such that $|f^{\prime}(x)|\geq\delta_{0}$ for all $x\in[\displaystyle\inf_{n}a_{n},\infty)$. Then, for almost all $\alpha\in[0,1]$ and for all sufficiently large $N$, we have
$\delta_{\mathrm{min}}^{\alpha}((f(a_{n}))_{n\geq 1},N)\geq\frac{1}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}((f(a_{n}))_{n\geq 1},N)\leq 1-\frac{1}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},$
where $\widetilde{D}_{N}=\#(f(A_{N})-f(A_{N}))^{+}$ and $A_{N}=\{a_{1},a_{2},\dots,a_{N}\}$.
Proof.
By the mean value theorem, for any $m\neq n$, there exists $\xi\in[\inf(a_{m},a_{n}),\sup(a_{m},a_{n})]$ such that
$|f(a_{m})-f(a_{n})|=|f^{\prime}(\xi)||a_{m}-a_{n}|\geq\delta_{0}|a_{m}-a_{n}|.$
Therefore, for all $m\neq n$,
(42) $|f(a_{m})-f(a_{n})|\geq\delta_{0}|a_{m}-a_{n}|\geq\delta_{0}\inf_{k\neq\ell}|a_{k}-a_{\ell}|\geq c\delta_{0}.$
Hence, $\inf_{m\neq n}|f(a_{m})-f(a_{n})|\geq c\delta_{0}>0$.
Since $|f^{\prime}(x)|\geq\delta_{0}>0$ for all $x$ and $f$ is differentiable, $f$ is injective. Thus, $(f(a_{n}))_{n\geq 1}$ is a sequence of distinct positive real numbers satisfying $\inf_{m\neq n}|f(a_{m})-f(a_{n})|\geq c\delta_{0}>0$. We obtain the required bounds by applying Theorem 1. ∎
Corollary 22.
Let $\epsilon>0$. Let $f$ be a differentiable function on $1,\infty)$, which is positive-valued on $\mathbb{N}$, and for which there exists $\delta_{0}>0$ such that $|f^{\prime}(x)|\geq\delta_{0}$ for all $x\in[1,\infty)$. Then, for almost all $\alpha\in[0,1]$ and for all sufficiently large $N$, we have
$\delta_{\mathrm{min}}^{\alpha}((f(n))_{n\geq 1},N)\geq\frac{1}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}((f(n))_{n\geq 1},N)\leq 1-\frac{1}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},$
where $\widetilde{D}_{N}=\#(f(A_{N})-f(A_{N}))^{+}$ and $A_{N}=\{1,2,\dots,N\}$.
Proof.
For $a_{n}=n$, the set $\{|a_{m}-a_{n}|:\,m\neq n\in\mathbb{N}\}$ is bounded below ($1$ is a lower bound) and has an infimum that satisfies $\inf_{m\neq n}|a_{m}-a_{n}|\geq c$ with $c=1$. The proof now follows from Corollary 21. ∎
We now prove Theorem 5, a variant of Theorem 1 in which the constant $c$ is replaced by a monotonically decreasing sequence $(c(N))_{N\geq 1}$.
3.2. Proof of Theorem 5
We first arrange the elements of $A=\displaystyle\bigcup_{N\geq 1}\left(A_{N}-A_{N}\right)^{+}$ in a sequence $\left(z_{n}\right)_{n\geq 1}$. To ensure a key monotonicity property, we choose this enumeration carefully.
For each $T\geq 2$, define the set of new differences in step $T$ by
(43) $B_{T}=(A_{T}-A_{T})^{+}\setminus(A_{T-1}-A_{T-1})^{+}.$
We now define the sequence $(z_{n})_{n\geq 1}$ by first listing all the elements of $B_{2}$ (in any order), then all the elements of $B_{3}$, and so on. This construction ensures that for all $N$, we have
(44) $\left\{z_{n}:1\leq n\leq D_{N}\right\}=\left(A_{N}-A_{N}\right)^{+},$
with $(0=)D_{1}<D_{2}<\ldots<D_{N}<D_{N+1}<\ldots$. Observe that $D_{N}$ satisfies $D_{N}\geq N-1$.
For each $n\in\mathbb{N}$, we define $f(n)$ as the smallest integer $T$ such that $z_{n}\in(A_{T}-A_{T})^{+}$. By our construction, if $z_{n}\in B_{T}$, then $f(n)=T$. Since the sequence $(z_{n})_{n\geq 1}$ is ordered by the sets $B_{2},B_{3},\ldots$, the function $n\mapsto f(n)$ is non-decreasing. Also, $z_{n}\in(A_{f(n)}-A_{f(n)})^{+}$ and therefore
(45) $z_{n}\geq\inf(A_{f(n)}-A_{f(n)})^{+}=\inf_{\begin{subarray}{c}1\leq r\neq s\leq f(n)\end{subarray}}|a_{r}-a_{s}|\geq c(f(n)).$
For each $n\in\mathbb{N}$, we set
(46) $\psi(n)=\frac{c(f(n))}{n(\log\sqrt{n})(\log_{2}\sqrt{n})^{1+\epsilon}}.$
Note that $(f(n))_{n\geq 1}$ is non-decreasing and $(c(n))_{n\geq 1}$ is decreasing, implying that $(c(f(n)))_{n\geq 1}$ is decreasing. Consequently, $(\psi(n))_{n\geq 1}$ is also decreasing.
For each $n\in\mathbb{N}$, we define
(47) $S_{n}=[0,1]\cap\bigcup_{a=0}^{\lfloor z_{n}\rfloor+1}\left(\frac{a}{z_{n}}-\frac{\psi(n)}{z_{n}},\frac{a}{z_{n}}+\frac{\psi(n)}{z_{n}}\right).$
Using (45), we see that the Lebesgue measure of $S_{n}$ satisfies
(48) $\lambda(S_{n})\leq(\lfloor z_{n}\rfloor+2)\frac{2\psi(n)}{z_{n}}\leq 2\psi(n)\left(1+\frac{2}{z_{n}}\right)\leq 2\psi(n)\left(1+\frac{2}{c(f(n))}\right).$
Substituting $\psi(n)$, we obtain
(49) $\lambda(S_{n})\leq\frac{2c(f(n))+4}{n(\log\sqrt{n})(\log_{2}\sqrt{n})^{1+\epsilon}}.$
The sequence $(c(f(n)))_{n\geq 1}$ being a monotone decreasing sequence of positive reals has an upper bound, namely $c(f(1))$. Therefore, summing over all $n\in\mathbb{N}$, we obtain
(50) $\sum_{n=1}^{\infty}\lambda(S_{n})\leq\sum_{n=1}^{\infty}\frac{2c(f(1))+4}{n(\log\sqrt{n})(\log_{2}\sqrt{n})^{1+\epsilon}}<\infty,$
where convergence follows from the Cauchy condensation test. Therefore, by the Borel-Cantelli lemma, there exists a set $\tilde{Q}\subseteq[0,1]$ with $\lambda(\tilde{Q})=1$ (without loss of generality, we can assume that $\tilde{Q}\cap\mathbb{Q}=\phi$) such that for all $\alpha\in\tilde{Q}$ and all $n\geq N_{1}(\alpha)$, we have $\alpha\notin S_{n}$, that is,
(51) $|\alpha z_{n}-a|\geq\psi(n),\text{ for all }a\in\{0,1,\ldots,\lfloor z_{n}\rfloor+1\}.$
Then, for all $\alpha\in[0,1]$, we have $0\leq\alpha z_{n}\leq z_{n}<\lfloor z_{n}\rfloor+1$. Therefore, (51) ensures that for all $\alpha\in\tilde{Q}$ and for all $n\geq N_{1}(\alpha)$, we have
[…]
(a) Let $a_{m},a_{n}\in A_{N}$ with $m\neq n$. Then, $|a_{m}-a_{n}|=z_{\tilde{n}}$ for some $\tilde{n}(=\tilde{n}(m,n))\leq D_{N}$, where $D_{N}\leq N^{2}$. $f(D_{N})$ is the smallest $T$ such that $z_{D_{N}}\in(A_{T}-A_{T})^{+}$ and $z_{D_{N}}\in(A_{N}-A_{N})^{+}$. Therefore, $f(D_{N})\leq N$ for all $N$, and since $(c(n))_{n\geq 1}$ is decreasing, we have $c(f(D_{N}))\geq c(N)$ for all $N$.
In the next few paragraphs, we prove $\|\alpha z_{\tilde{n}}\|>\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$ for all sufficiently large $N$. We consider a finite set of indices $\{1,2,\dots,N_{1}(\alpha)-1\}$. Since $z_{n}$ are positive and distinct, and since $\alpha$ is such that $\alpha z_{n}$ is never an integer for any $n$ (which holds for almost all $\alpha$), there exists a constant $\eta(\alpha)>0$ such that
(53) $\|\alpha z_{n}\|\geq\eta(\alpha)\quad\text{for all}\quad n<N_{1}(\alpha).$
Furthermore, $D_{N}\geq N-1$ implies that the sequence
(54) $\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\to 0\quad\text{as}\quad N\to\infty.$
Therefore, there exists $N_{2}(\alpha)$ such that for all $N\geq N_{2}(\alpha)$,
(55) $\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}<\eta(\alpha).$
Now, let $N_{0}(\alpha)=\max\{N_{1}(\alpha),N_{2}(\alpha)\}$. Then, for all $N\geq N_{0}(\alpha)$ and for every pair of distinct elements $a_{m},a_{n}\in A_{N}$ with the corresponding index $\tilde{n}$:
  • If $\tilde{n}\geq N_{1}(\alpha)$, then using the facts that $\psi$ is decreasing and $\tilde{n}\leq D_{N}$, we have
    (56) $\|\alpha z_{\tilde{n}}\|\underbrace{\geq}_{\tilde{n}\geq N_{1}(\alpha)}\psi(\tilde{n})\geq\psi(D_{N})=\frac{c(f(D_{N}))}{D_{N}(\log\sqrt{D_{N}})(\log_{2}\sqrt{D_{N}})^{1+\epsilon}}\geq\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}.$
  • If $\tilde{n}<N_{1}(\alpha)$, then
    (57) $\|\alpha z_{\tilde{n}}\|\underbrace{\geq}_{\tilde{n}<N_{1}(\alpha)}\eta(\alpha)\underbrace{>}_{N\geq N_{0}(\alpha)}\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}.$
Therefore, for all $N\geq N_{0}(\alpha)$, and for all distinct $a_{m},a_{n}\in A_{N}$, we have
[…]
for almost all $\alpha\in[0,1]$. The proof now follows by taking the minimum over all $a_{m},a_{n}\in A_{N}$.
(b) Since $0\leq\alpha z_{n}\leq z_{n}$, we have $0\leq\lfloor\alpha z_{n}\rfloor\leq\lfloor z_{n}\rfloor.$ In particular, taking $a=\lfloor\alpha z_{n}\rfloor$ in (51), we obtain that for all $\alpha\in\tilde{Q}$ (with $\lambda(\tilde{Q})=1$) and for all $n\geq N_{0}(\alpha)$, we have
(59) $\{\alpha z_{n}\}\geq\psi(n).$
Therefore, for all $N\geq N_{0}(\alpha)$ and for all $a_{m},a_{n}\in A_{N}$ with ${m}\neq{n}$, the inequality
(60) $\left\{\alpha\left|a_{m}-a_{n}\right|\right\}\geq\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$
holds for almost all $\alpha\in[0,1]$.
Note that t= $t||=\text{ min }(\{t\},1-\{t\})\leq 1-\{$t\} and t=t||-t||=||t|| for any $t\in\mathbb{R}$. Therefore, using (59), for all $N\geq N_{0}(\alpha)$, and for all distinct $a_{m},a_{n}\in A_{N}$, we have
(61) $\left\|\alpha\left(a_{m}-a_{n}\right)\right\|=\left\|\alpha|a_{m}-a_{n}|\right\|\leq 1-\{\alpha|a_{m}-a_{n}|\}\leq 1-\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$
for almost all $\alpha\in[0,1]$. The proof now follows by taking the minimum over all $a_{m},a_{n}\in A_{N}$.
Corollary 23.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq c$, where $c$ is a positive constant. Let $\epsilon>0$, and let $D_{N}$ be defined as in Section 2.1. Then, for almost all $\alpha\in[0,1]$, there exists $N_{0}(\alpha)$ such that for all $N\geq N_{0}(\alpha)$, we have
(62) $\delta^{\alpha}_{\min}(N)\geq\dfrac{\max(1,c)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}.$
Proof.
Taking $c(N)=c$ for all $N$, in Theorem 5, we obtain that for any $\epsilon>0$ and for almost all $\alpha\in[0,1]$, $\delta^{\alpha}_{\min}(N)\geq\dfrac{c}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$ for all sufficiently large $N$. The proof now follows by combining this observation with Theorem 1. ∎
Remark 24.
In particular, the Fibonacci sequence $(F_{n})_{n\geq 2}$ (with $F_{2}=1,F_{3}=2,F_{4}=3,\dots$) satisfies $\inf_{m\neq n}|F_{m}-F_{n}|=1$, and hence is a suitable choice for $a_{n}$ in Corollary 23.
Remark 25 (Beatty sequence).
[…]
$\delta_{\mathrm{min}}^{\alpha}((\lfloor n\theta\rfloor)_{n\geq 1},N)\geq\dfrac{\lfloor\theta\rfloor}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\geq\dfrac{\lfloor\theta\rfloor}{N^{2}\log N(\log_{2}N)^{1+\epsilon}}.$
In particular, one can consider the sequences $\lfloor\sqrt{2}n\rfloor$ and $\lfloor n\phi\rfloor$, where $\phi=\frac{1+\sqrt{5}}{2}$ is the golden ratio.
Example 26.
For positive constants $a$ and $d$, let $a_{n}=a+(n-1)d$. Then $D_{N}=\#(A_{N}-A_{N})^{+}=\#\{d,2d,\ldots,(N-1)d\}=N-1\leq N$, and we obtain
$\delta^{\alpha}_{\min}(N)\geq\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\geq\frac{1}{N\log N(\log_{2}N)^{1+\epsilon}},$
which is $\gg\frac{1}{N(\log N)^{1+\epsilon}}$ for any $\epsilon>0$.
In particular, taking $a=d=1$, that is, for the sequence $a_{n}=n$, we recover the result obtained in [Rudnick, Section 3].
Example 27.
Let $\theta\geq 1$ and $a_{n}=n^{\theta}$. Since $(a_{n})_{n\geq 1}$ is a strictly increasing sequence and $f(x)=(x+1)^{\theta}-x^{\theta}$ is a monotone increasing function in $1,\infty)$, we have $|a_{m}-a_{n}|\geq|a_{n+1}-a_{n}|=(n+1)^{\theta}-n^{\theta}\geq 2^{\theta}-1$ for all $m>n$. This ensures that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|=\displaystyle\inf_{\begin{subarray}{c}1\leq n<m\leq N\end{subarray}}|a_{m}-a_{n}|\geq 2^{\theta}-1$. Therefore, using Corollary 23, we obtain that for almost all $\alpha\in[0,1]$, and for all sufficiently large $N$,
$\delta_{\mathrm{min}}^{\alpha}((n^{\theta})_{n\geq 1},N)\geq\frac{2^{\theta}-1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\geq\frac{2^{\theta}-1}{N^{2}\log N(\log_{2}N)^{1+\epsilon}},\text{ and }$
$\delta_{\mathrm{min}}^{\alpha}((n^{\theta})_{n\geq 1},N)\leq 1-\frac{2^{\theta}-1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\leq 1-\frac{2^{\theta}-1}{N^{2}\log N(\log_{2}N)^{1+\epsilon}},\text{ where}$
$A_{N}=\{1^{\theta},2^{\theta},\dots,N^{\theta}\}$ and $D_{N}=\#(A_{N}-A_{N})^{+}$.
Remark 28.
The following functions can be taken in Corollary 22 as suitable choices for $f(x)\colon$ $\sinh(x)=\frac{e^{x}-e^{-x}}{2}$, $e^{x}$, $e^{x}+x^{k}$, $x^{k}\log{(x+1)}$, $x\log x$, where $k\in\mathbb{N}$.
Example 29.
A sequence $(a_{n})_{n\geq 1}$ of positive real numbers is called lacunary if there exists a constant $D>1$ such that $\frac{a_{n+1}}{a_{n}}\geq D$ for all $n\in\mathbb{N}$.
A lacunary sequence satisfies all the conditions of Theorem 1.
  1. (1)
    It consists of distinct positive real numbers (since $a_{n+1}\geq Da_{n}>a_{n}$ for all $n$).
  2. (2)
    For $m\neq n\in\mathbb{N}$, using lacunarity, we obtain
    (63) $|a_{m}-a_{n}|\geq a_{\max(m,n)}-a_{\max(m,n)-1}\geq(D-1)a_{\max(m,n)-1}\geq(D-1)a_{1}.$
    That is, $(a_{n})_{n\geq 1}$ satisfies the condition $\inf_{m\neq n}|a_{m}-a_{n}|\geq c$ with $c=(D-1)a_{1}>0$.
Therefore, Theorem 1 applies to lacunary sequences. Since the sequence $a_{n}\coloneqq c^{n}$, where $c>1$ is a fixed constant, is lacunary, Theorem 1 also applies to such sequences, in particular, to the sequences $\left(e^{n}\right)_{n\geq 1}$, $\left(\pi^{n}\right)_{n\geq 1}$, and $\left((\frac{3}{2})^{n}\right)_{n\geq 1}$.
Remark 30.
Let $f$ satisfy the hypotheses in Corollary 21, and let $(a_{n})_{n\geq 1}$ be a lacunary sequence. Then, for almost all $\alpha\in[0,1]$ and for all sufficiently large $N$, we have
$\delta_{\mathrm{min}}^{\alpha}((f(a_{n}))_{n\geq 1},N)\geq\frac{1}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}((f(a_{n}))_{n\geq 1},N)\leq 1-\frac{1}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},$
where $\widetilde{D}_{N}=\#(f(A_{N})-f(A_{N}))^{+}$.
Corollary 31.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq c(N)$, where $(c(N))_{N\geq 1}$ is a monotone decreasing sequence of positive reals. Let $f$ be a differentiable function on $\inf_{n}a_{n},\infty)$, which is positive-valued on $(a_{n})_{n\geq 1}$ and for which there exists $\delta_{0}>0$ such that $|f^{\prime}(x)|\geq\delta_{0}$ for all $x\in[\displaystyle\inf_{n}a_{n},\infty)$. Then, for almost all $\alpha\in[0,1]$ and for all sufficiently large $N$, we have
$\delta_{\mathrm{min}}^{\alpha}((f(a_{n}))_{n\geq 1},N)\geq\frac{c(N)\delta_{0}}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}((f(a_{n}))_{n\geq 1},N)\leq 1-\frac{c(N)\delta_{0}}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},$
where $\widetilde{D}_{N}=\#(f(A_{N})-f(A_{N}))^{+}$.
Proof.
By the mean value theorem, for any $m\neq n$, there exists $\xi\in[\inf(a_{m},a_{n}),\sup(a_{m},a_{n})]$ such that
$|f(a_{m})-f(a_{n})|=|f^{\prime}(\xi)||a_{m}-a_{n}|\geq\delta_{0}|a_{m}-a_{n}|.$
Therefore, for all $1\leq m^{\prime},n^{\prime}\leq N$ with $m^{\prime}\neq n^{\prime}$,
$|f(a_{m^{\prime}})-f(a_{n^{\prime}})|\geq\delta_{0}|a_{m^{\prime}}-a_{n^{\prime}}|\geq\delta_{0}\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq c(N)\delta_{0}.$
In particular, $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|f(a_{m})-f(a_{n})|\geq d(N)$, where $d(N)=c(N)\delta_{0}$.
Since $c(N)$ is a monotone decreasing sequence of positive reals, so is $d(N)$. Since $|f^{\prime}(x)|\geq\delta_{0}>0$ for all $x$ and $f$ is differentiable, $f$ is injective. Thus, $(f(a_{n}))_{n\geq 1}$ is a sequence of distinct positive real numbers satisfying $\inf_{1\leq m\neq n\leq N}|f(a_{m})-f(a_{n})|\geq d(N)$. We obtain the required bound in part (a) by applying Theorem 5. The proof of part (b) follows from Theorem 5. ∎
Example 32.
Let $\theta<1$ and $f(x)=x^{\theta}$. Then $f$ is differentiable and positive-valued on $\mathbb{N}$, satisfying $|f^{\prime}(x)|=|\theta||x|^{\theta-1}\geq|\theta|N^{\theta-1}$ for all $x\in[1,N]$. Let $a_{n}=f(n)=n^{\theta}$ for all $n\in\mathbb{N}$. Therefore, by the mean value theorem, $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq\displaystyle\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|\theta|N^{\theta-1}|m-n|\geq|\theta|N^{\theta-1}$. The assumption $\theta<1$ ensures that the sequence $(c(N))_{N\geq 1}$, where $c(N):=|\theta|N^{\theta-1}$ is a monotone decreasing.
Therefore, using Theorem 5, we obtain that for almost all $\alpha\in[0,1]$ and for all sufficiently large $N$,
$\delta_{\mathrm{min}}^{\alpha}((n^{\theta})_{n\geq 1},N)\geq\frac{|\theta|N^{\theta-1}}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\geq\frac{|\theta|}{N^{3-\theta}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}((n^{\theta})_{n\geq 1},N)\leq 1-\frac{|\theta|N^{\theta-1}}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}\leq 1-\frac{|\theta|}{N^{3-\theta}\log N(\log_{2}N)^{1+\epsilon}},\text{ where}$
$A_{N}=\{1^{\theta},2^{\theta},\dots,N^{\theta}\}$ and $D_{N}=\#(A_{N}-A_{N})^{+}$.
In particular, when $\theta=\frac{1}{2}$, we have
$\delta_{\mathrm{min}}^{\alpha}((\sqrt{n})_{n\geq 1},N)\geq\frac{1}{2N^{5/2}\log N(\log_{2}N)^{1+\epsilon}}$
for almost all $\alpha\in[0,1]$ and for all sufficiently large $N$. We remark that the lower bound we obtained for the minimal gap of the sequence $(\sqrt{n})_{n\geq 1}$ is weaker than the optimal bound, which is $\sim\frac{1}{2N^{3/2}}$ ((see [Regavim, Theorem $1.9$])). For any $\theta$ which is very close to 11 $(\theta<1)$, the lower bound we obtain for the minimal gap of the sequence $(n^{\theta})_{n\geq 1}$ is nearly optimal, which is believed to be $\sim\frac{1}{N^{2}}$.
Corollary 33.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that
(64) $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq N^{\delta-1}$
for some $\delta\in[0,1)$. Let $f$ be a differentiable function on $\inf_{n}a_{n},\infty)$, which is positive-valued on $(a_{n})_{n\geq 1}$ and for which there exists $\delta_{0}>0$ such that $|f^{\prime}(x)|\geq\delta_{0}$ for all $x\in[\displaystyle\inf_{n}a_{n},\infty)$. Then, for almost all $\alpha\in[0,1]$ and for all sufficiently large $N$, we have
$\delta_{\mathrm{min}}^{\alpha}((f(a_{n}))_{n\geq 1},N)\geq\frac{N^{\delta-1}\delta_{0}}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}}\geq\frac{\delta_{0}}{N^{3-\delta}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}((f(a_{n}))_{n\geq 1},N)\leq 1-\frac{N^{\delta-1}\delta_{0}}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}}\leq 1-\frac{\delta_{0}}{N^{3-\delta}\log N(\log_{2}N)^{1+\epsilon}}.$
where $\widetilde{D}_{N}=\#(f(A_{N})-f(A_{N}))^{+}$.
Proof.
The proof follows immediately from Corollary 31 taking $c(N)=N^{\delta-1}$, where $\delta<1$. ∎
Example 34.
Let $f(x)=\log(1+x)$ and $a_{n}=f(n)$. Then, the mean value theorem ensures that there exists $c_{m,n}\in(\inf(m,n),\sup(m,n))$ such that $|a_{m}-a_{n}|=|f^{\prime}(c_{m,n})||m-n|\geq\frac{1}{1+c_{m,n}}$, that is, $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq\frac{1}{N+1}\geq\frac{1}{2N}$. Therefore, for any $\epsilon>0$ and for almost all $\alpha\in[0,1]$, there exists $N_{0}(\alpha)$ such that for all $N\geq N_{0}(\alpha)$, $\delta^{\alpha}_{\min}((\log(1+n))_{n\geq 1},N)\geq\dfrac{1}{2D_{N}N\log N(\log_{2}N)^{1+\epsilon}}\geq\dfrac{1}{2N^{3}\log N(\log_{2}N)^{1+\epsilon}}$, where $A_{N}=\{\log 2,\log 3,\dots,\log(1+N)\}$ and $D_{N}=\#(A_{N}-A_{N})^{+}$.
4. Proofs of Theorems 69 and 11: Upper Bounds
The convergence part of the Borel–Cantelli lemma holds unconditionally, whereas the divergence part requires the hypothesis of pairwise independence. Consequently, proving divergence in the context of metric number theory often involves greater technical difficulty than proving convergence (see [BV2023] for a discussion). In Section 4.1, we briefly outline the heuristics behind the proof. In Section 4.2, we provide a proof of Theorem 6, where the upper bound depends on the size of $a_{N}$. Section 4.3.1 compiles the necessary technical lemmas and overlap estimates required for the subsequent analysis in Theorem 9, and in Section 4.3.2, we finally present the proof of Theorem 9, where the upper bound is independent of the size of $a_{N}$. In Section 4.4, we prove Theorem 11, where the upper bound result holds for all except finitely many $N$.
4.1. Outline and heuristics
The proofs follow the modern approach to metric Diophantine approximation problems. The fundamental challenge is that the approximation sets under consideration lack the stochastic independence required for a direct application of the second Borel-Cantelli lemma, as demonstrated by the classical Duffin-Schaeffer counterexample (see [Duffin1941KhintchinesPI]).
Instead, we rely on the concept of “quasi-independence on average” (See Lemma 119 and [BDV2006, BV2010, BV2023]), where the idea is to define the $S_{n}$ such that the sum of their measures diverges, while simultaneously controlling the measure of their pairwise overlaps ($S_{m}\cap S_{n}$ for $m\neq n$) to apply a suitable convergence criterion. Following Rudnick’s strategy, we employ $L^{2}$ methods that naturally lead to the study of greatest common divisor (GCD) sums, which play a central role in this area.
Recent optimal bounds for GCD sums obtained by de la Bretèche and Tenenbaum (see [delabretèche]) introduce a specific growth factor that determines the threshold for successful application of this method. Specifically, for the $L^{2}$ method to succeed, the sum of measures of our approximation sets must exceed this GCD sum growth factor. This explains the precise form of the conditions appearing in our theorems.
4.2. Proof of Theorem 6
We begin by stating a conjecture that is now a theorem due to the work of Koukoulopoulos and Maynard (see [KoukouMaynard]). This result will be useful in the proof of Theorem 6.
Theorem 35 (Duffin-Schaeffer Conjecture).
Let $\psi:\mathbb{N}\to\mathbb{R}_{\geq 0}$ be a function such that $\sum_{q=1}^{\infty}\frac{\psi(q)\phi(q)}{q}=\infty$. Let $\mathcal{A}$ be the set of $\alpha\in[0,1]$ for which the inequality
(65) $\left|\alpha-\frac{a}{q}\right|\leq\frac{\psi(q)}{q},\,(a,q)=1,\,q\geq 1$
has infinitely many solutions in $a$ and $q$. Then $\mathcal{A}$ has the Lebesgue measure $1$.
Remark 36.
Here we restrict our attention to $\alpha\in[0,1]$ only, because if $\alpha$ has an approximation given by inequality (65), then so does $\alpha+r$ for any integer $r$, with $a$ replaced by $a+rq$. Note that $(a,q)=1$ implies $(a+rq,q)=1$.
Lemma 37.
Let $(a_{n})_{n\geq 1}$ be a sequence of real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Define the finite set $A_{N}=\{a_{1},\dots,a_{N}\}$ and let $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$. Then $H_{N}$ satisfies the recurrence $\lceil\frac{N}{2}\rceil\leq H_{N+1}\leq H_{N}+N$. Consequently, $H_{N+1}\leq 3H_{N}+1\leq 4H_{N}$ holds for $N\geq 3$.
Proof.
We begin by expressing the positive difference set for $A_{N+1}$ as
(66) $(A_{N+1}-A_{N+1})^{+}=(A_{N}-A_{N})^{+}\cup\{|a_{N+1}-a|:a\in A_{N}\}.$
Taking the floor of both sides (applied element-wise to the set) gives
(67) $\lfloor(A_{N+1}-A_{N+1})^{+}\rfloor=\lfloor(A_{N}-A_{N})^{+}\rfloor\cup S,$
where $S=\{\lfloor|a_{N+1}-a|\rfloor:a\in A_{N}\}$. It follows that
(68) $H_{N+1}=\#\lfloor(A_{N+1}-A_{N+1})^{+}\rfloor\leq\#\lfloor(A_{N}-A_{N})^{+}\rfloor+\#S\leq H_{N}+N,$
which establishes the upper bound.
For the lower bound, consider the set of floor values of distances from $a_{1}$:
(69) $B=\{\lfloor|a_{j}-a_{1}|\rfloor:2\leq j\leq N\}.$
Each element in $B$ is an integer and we claim that for each integer $n\in\mathbb{N}$, there can be at most two indices $j\geq 2$ such that $\lfloor|a_{j}-a_{1}|\rfloor=n$.
Suppose for contradiction that there exist three distinct indices $j,k,\ell\geq 2$ with
(70) $\lfloor|a_{j}-a_{1}|\rfloor=\lfloor|a_{k}-a_{1}|\rfloor=\lfloor|a_{\ell}-a_{1}|\rfloor=n.$
This implies $n\leq|a_{j}-a_{1}|,|a_{k}-a_{1}|,|a_{\ell}-a_{1}|<n+1$. By the pigeonhole principle, at least two of these points must be on the same side of $a_{1}$ (that is, both larger or both smaller than $a_{1}$). Without loss of generality, assume that both $a_{j}$ and $a_{k}$ are larger than $a_{1}$ (the case where both are smaller is similar). Then $a_{j},a_{k}\in[a_{1}+n,a_{1}+n+1)$.
But then $|a_{j}-a_{k}|<1$, contradicting the assumption that $\inf_{m\neq n}|a_{m}-a_{n}|\geq 1$. Hence, for each $n\in\mathbb{N}$, there are at most two points $a_{j}$ (with $j\geq 2$) such that $\lfloor|a_{j}-a_{1}|\rfloor=n$.
Therefore, the number of distinct values in $B$ is at least $\left\lceil\frac{N-1}{2}\right\rceil$. Since $B\subseteq\lfloor|A_{N}-A_{N}|\rfloor$, it follows that the number of distinct distances $H_{N}$ satisfies $H_{N}\geq\left\lceil\frac{N-1}{2}\right\rceil$. This proves the first part of the lemma.
For $N\geq 3$, the lower bound gives $H_{N}\geq\frac{N-1}{2}(\geq 1)$. Substituting this into the upper bound gives
(71) $H_{N+1}\leq H_{N}+2\left(\frac{N-1}{2}\right)+1\leq H_{N}+2H_{N}+1\leq 4H_{N}.$
This completes the proof. ∎
Remark 38.
Lemma 37 shows that $H_{N+1}\leq H_{N}+N$, and by iterating this recurrence, we obtain the trivial bound $H_{N}\leq\displaystyle\sum_{k=1}^{N-1}k=\binom{N}{2}$.
We now establish a variant of Lemma 37 for strictly increasing sequences, which yields an improved bound for $H_{N}$. Although this lemma is not necessary for the proof of Theorem 6, it will be used in the proof of one of its corollaries.
Lemma 39.
Let $(a_{n})_{n\geq 1}$ be a strictly increasing sequence of real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Then $H_{N}$ satisfies the recurrence $N\leq H_{N+1}\leq H_{N}+N$. Consequently, $H_{N+1}\leq 2H_{N}+1\leq 3H_{N}$ holds for $N\geq 2$.
Proof.
We express the floor of positive difference set for $A_{N+1}$ as
(72) $\lfloor(A_{N+1}-A_{N+1})^{+}\rfloor=\lfloor(A_{N}-A_{N})^{+}\rfloor\cup S,$
where $S=\{\lfloor a_{N+1}-a\rfloor:a\in A_{N}\}$. We now claim that the set $S$ contains $N$ distinct elements.
Suppose, for contradiction, that this is not true. Then, there exist indices $1\leq i\neq j\leq N$ such that
(73) $\lfloor a_{N+1}-a_{i}\rfloor=\lfloor a_{N+1}-a_{j}\rfloor=k\quad\text{(say)}.$
By the definition of the floor function, we can write $a_{N+1}-a_{i}=k+t_{i}$ and $a_{N+1}-a_{j}=k+t_{j}$, where $0\leq t_{i},t_{j}<1$. Therefore,
(74) $|a_{j}-a_{i}|=|(a_{N+1}-a_{i})-(a_{N+1}-a_{j})|=|(k+t_{i})-(k+t_{j})|=|t_{i}-t_{j}|\in[0,1).$
This implies $\lfloor|a_{i}-a_{j}|\rfloor=0$ for $i\neq j$, which contradicts the initial assumption that $\inf_{m\neq n}|a_{m}-a_{n}|\geq 1$. This contradiction proves that $S$ has $N$ distinct elements. Moreover, since the sequence is strictly increasing, all elements of $S$ are positive, and we have $H_{N+1}\geq\#S=N$.
The upper bound follows directly from Lemma 37. This proves the first part of the lemma.
For $N\geq 2$, the lower bound gives $H_{N}\geq N-1(\geq 1)$. Substituting this into the upper bound gives
(75) $H_{N+1}\leq H_{N}+(N-1)+1\leq 2H_{N}+1\leq 3H_{N}.$
This completes the proof. ∎
Remark 40.
Alternatively, to show $H_{N}\geq N-1$ for an increasing sequence $(a_{n})_{n\geq 1}$ with $\inf_{m\neq n}|a_{m}-a_{n}|\geq 1$, consider the set $\mathcal{A}\coloneqq\{\lfloor a_{j}-a_{1}\rfloor:2\leq j\leq N\}$. Since $a_{j}-a_{1}=(a_{j}-a_{j-1})+(a_{j-1}-a_{1})\geq 1+(a_{j-1}-a_{1})$, we have $\lfloor a_{j}-a_{1}\rfloor\geq\lfloor a_{j-1}-a_{1}\rfloor+1$. Therefore, the values $\lfloor a_{j}-a_{1}\rfloor$ increase strictly with $j$, and therefore are all distinct. Since $\mathcal{A}\subseteq\lfloor(A_{N}-A_{N})^{+}\rfloor$, it follows that $H_{N}\geq N-1$.
Proof of Theorem 6. We arrange the elements of $A=\displaystyle\cup_{N\geq 1}\lfloor(A_{N}-A_{N})^{+}\rfloor$ into a sequence. Thus, let $(Z_{n})_{n\geq 1}$ be a sequence of distinct positive integers such that we have
(76) $\{Z_{n}:1\leq n\leq H_{N}\}=\lfloor(A_{N}-A_{N})^{+}\rfloor\text{ for all }N.$
The sequence $(Z_{n})_{n\geq 1}$ is not uniquely defined, as $H_{N+1}-H_{N}$ can be strictly larger than $1$. In fact, it can be as large as $N$, but the proof works irrespective of the version of $(Z_{n})_{n\geq 1}$ we take as long as (76) is satisfied. To reduce the size of overlaps $S_{m}\cap S_{n}$, we use the same strategy as is used in the coprime setup of the Duffin-Schaeffer conjecture. We replace $S_{m}$ and $S_{n}$ with modified sets that preserve most of the measure of the original sets, but remove those parts that are excessively responsible for the overlaps. To incorporate this idea into our situation, we consider the reduced sets
(77) $S_{n}^{\text{coprime }}=[0,1]\cap\left(\bigcup_{\begin{subarray}{c}0\leq a\leq Z_{n}\\ \gcd\left(a,Z_{n}\right)=1\end{subarray}}\left(\frac{a}{Z_{n}}-\frac{\psi(n)}{Z_{n}},\frac{a}{Z_{n}}+\frac{\psi(n)}{Z_{n}}\right)\right),$
where $\psi(n)=\frac{\log_{2}Z_{n}}{n\log n\log_{2}n}$ for all $n$. This reflects the fact that the overlaps $S_{m}\cap S_{n}$ come mainly from intervals in $S_{m}$ and $S_{n}$, which are centred around points $a/Z_{m}$ and $b/Z_{n}$, respectively, for which $a/Z_{m}=$ $b/Z_{n}$, so that either $\gcd\left(a,Z_{m}\right)>1$ or $\gcd\left(b,Z_{n}\right)>1$.
Let
(78) $S_{n}=[0,1]\cap\left(\bigcup_{\begin{subarray}{c}0\leq a\leq Z_{n}\end{subarray}}\left(\frac{a}{Z_{n}}-\frac{\psi(n)}{Z_{n}},\frac{a}{Z_{n}}+\frac{\psi(n)}{Z_{n}}\right)\right).$
Then
$\lambda(S_{n})=\frac{2\psi(n)}{Z_{n}}Z_{n}=\frac{2\log_{2}Z_{n}}{n\log n\log_{2}n}.$
It is known that ${\varphi(T)/T}\gg{(\log_{2}T)^{-1}}$, that is, there exist a positive constant $C_{0}$ and an $N_{0}^{(1)}\in\mathbb{N}$ such that ${\varphi(Z_{n})/Z_{n}}\geq{C_{0}(\log_{2}Z_{n})^{-1}}$ for all $n\geq N_{0}^{(1)}$. Observe that this is a function of $Z_{n}$, which means that it depends on the actual size of the elements of the floored positive difference set $\lfloor(A_{N}-A_{N})^{+}\rfloor$, and not a function of the cardinality of the difference set, that is, $n$.
We first assume that $\psi(n)\geq\frac{1}{2}$ for infinitely many $n$. Since $\|\alpha z_{n}\|\leq\frac{1}{2}$, we have $\|\alpha z_{n}\|\leq\psi(n)$ for infinitely many $n$. Otherwise, $\psi(n)\geq\frac{1}{2}$ for only finitely many $n$, that is, there exists $N_{0}^{(2)}$ such that $\psi(n)<\frac{1}{2}$ for all $n\geq N_{0}^{(2)}$, which ensures that $2\psi(n)\varphi\left(Z_{n}\right)/Z_{n}<1$. Let $N_{0}\coloneqq\max(N_{0}^{(1)},N_{0}^{(2)})$. Therefore,
(79) $\sum_{n=N_{0}}^{\infty}\lambda\left(S_{n}^{\text{coprime }}\right)=$ $\sum_{n=N_{0}}^{\infty}\min(2\psi(n)\varphi\left(Z_{n}\right)/Z_{n},1)$
(80) $=$ $\sum_{n=N_{0}}^{\infty}2\psi(n)\varphi\left(Z_{n}\right)/Z_{n}$
(81) $\geq$ $\sum_{n=N_{0}}^{\infty}\frac{2C_{0}\psi(n)}{\log_{2}Z_{n}}=\sum_{n=N_{0}}^{\infty}\frac{2C_{0}}{n\log n\log_{2}n}=+\infty.$
Let $\psi^{*}:\mathbb{N}\to\mathbb{R}_{\geq 0}$ be defined by
$\psi^{*}(q)=\begin{cases}\psi(n)&\quad\text{if }q=Z_{n}\text{ for some }n\in\mathbb{N},\\ 0&\quad\text{otherwise.}\end{cases}$
Since $(Z_{n})_{n\geq 1}$ is a sequence of distinct positive integers, the map $\psi^{*}$ is well-defined. Therefore,
(82) $\sum_{q=1}^{\infty}\frac{\psi^{*}(q)\phi(q)}{q}=\sum_{n=1}^{\infty}\frac{\psi(n)\phi(Z_{n})}{Z_{n}}=\frac{1}{2}\sum_{n=N_{0}}^{\infty}\lambda\left(S_{n}^{\text{coprime }}\right)=\infty.$
We note that $\psi^{*}:\mathbb{N}\to\mathbb{R}_{\geq 0}$ is a function such that $\sum_{q=1}^{\infty}\frac{\psi^{*}(q)\phi(q)}{q}=\infty$. Therefore, using the Duffin-Schaeffer conjecture (see Theorem 35), there exists a set $\mathcal{K^{\prime}}\subseteq[0,1]$ with $\lambda(\mathcal{K^{\prime}})=1$ (without loss of generality, we can assume that $\mathcal{K^{\prime}}\cap\mathbb{Q}=\phi$) such that for all $\alpha\in\mathcal{K^{\prime}}$, the inequality
(83) $\left|\alpha-\frac{a}{q}\right|\leq\frac{\psi^{*}(q)}{q}$
has infinitely many coprime solutions $a$ and $q$.
We note that if $q\neq Z_{n}$ for some $n\in\mathbb{N}$, then $\psi^{*}(q)=0$, and (83) gives $|\alpha-\frac{a}{q}|=0$, that is, $\alpha=\frac{a}{q}\in\mathbb{Q}$, which is a contradiction. In other words, $q=Z_{n}$ for some $n\in\mathbb{N}$.
Therefore, for all $\alpha\in\mathcal{K}$, where $\mathcal{K}\subseteq[0,1]$ and $\lambda(\mathcal{K})=1$, the inequality
(84) $\left|\alpha-\frac{a}{Z_{n}}\right|\leq\frac{\psi^{*}(Z_{n})}{Z_{n}}$
(85) $\implies$ $\left|\alpha-\frac{a}{Z_{n}}\right|\leq\frac{\psi(n)}{Z_{n}}$
(86) $\implies$ $\left|\alpha Z_{n}-a\right|\leq\psi(n)$
(87) $\implies$ $\left\|\alpha Z_{n}\right\|\leq\psi(n)$
has infinitely many coprime solutions $a$ and $Z_{n}$, that is, $\left\|\alpha Z_{n}\right\|\leq\frac{\log_{2}Z_{n}}{n\log n\log_{2}n}$ for infinitely many $n$. This shows that there exists a sequence $((t(n))_{n\geq 1}$ of distinct natural numbers such that $\left\|\alpha Z_{t(n)}\right\|\leq\frac{\log_{2}Z_{t(n)}}{t(n)\log t(n)\log_{2}t(n)}$ for all $n\in\mathbb{N}$.
By the construction of $Z_{n}$, for every $n$, there is a unique $N$ such that $H_{N}<t(n)\leq H_{N+1}$, and Lemma 37 shows that $t(n)>H_{N}\geq\frac{1}{4}H_{N+1}\geq\frac{N}{8}$.
Since $B_{N+1}=\displaystyle\max_{1\leq m\neq n\leq N+1}\lfloor|a_{m}-a_{n}|\rfloor=\max\lfloor(A_{N+1}-A_{N+1})^{+}\rfloor=\max\{Z_{n}:1\leq n\leq H_{N+1}\}$, we have $Z_{t(n)}\leq B_{N+1}$. Therefore, for all $n\in\mathbb{N}$ with $t(n)\in(H_{N},H_{N+1}]$, we have
(88) $\left\|\alpha Z_{t(n)}\right\|\leq\frac{\log_{2}Z_{t(n)}}{t(n)\log t(n)\log_{2}t(n)}\leq\frac{\log_{2}B_{N+1}}{H_{N}\log(\frac{N}{8})\log_{2}(\frac{N}{8})}.$
In particular, there are infinitely many $N$ such that
(89) $\lfloor\delta^{\alpha}_{\min}\rfloor(N+1)=$ $\displaystyle\min_{1\leq m\neq n\leq N+1}\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|=\displaystyle\min_{l\leq H_{N+1}}\|\alpha Z_{l}\|\leq\|\alpha Z_{t(n)}\|\leq\frac{\log_{2}B_{N+1}}{H_{N}\log(\frac{N}{8})\log_{2}(\frac{N}{8})}.$
Using Lemma 37, we see that there are infinitely many $N$ such that
(90) $\lfloor\delta^{\alpha}_{\min}\rfloor(N)\leq\frac{\log_{2}B_{N}}{H_{N-1}\log(\frac{N-1}{8})\log_{2}(\frac{N-1}{8})}\leq\frac{4\log_{2}B_{N}}{H_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}.$
We now prove that $B_{N}\leq\lfloor\max(a_{1},\ldots,a_{N})\rfloor$.
For any $1\leq m\neq n\leq N$, we have $|a_{m}-a_{n}|\leq\max(a_{1},\ldots,a_{N})-\min(a_{1},\ldots,a_{N})$. Taking floors, we get
(91) $\lfloor|a_{m}-a_{n}|\rfloor\leq\lfloor\max(a_{1},\ldots,a_{N})-\min(a_{1},\ldots,a_{N})\rfloor.$
Since we have all $a_{i}>0$, we have $\min(a_{1},\ldots,a_{N})>0$, so $\lfloor|a_{m}-a_{n}|\rfloor\leq\lfloor\max(a_{1},\ldots,a_{N})\rfloor$, and we obtain the upper bound claimed for $B_{N}$.
Corollary 41.
Let $(a_{n})_{n\geq 1}$ be an increasing sequence of distinct positive real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Then, for almost all $\alpha\in[0,1]$, we have
(92) $\lfloor\delta^{\alpha}_{\min}\rfloor(N)\leq\frac{3\log_{2}B_{N}}{H_{N}\log(\frac{N}{4})\log_{2}(\frac{N}{4})}\text{ for infinitely many $N$,}$
where $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\displaystyle\min_{1\leq m\neq n\leq N}\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|$, $B_{N}=\displaystyle\max_{1\leq m\neq n\leq N}\lfloor|a_{m}-a_{n}|\rfloor$, and $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$. Since $B_{N}\leq\lfloor a_{N}\rfloor$, the above statement also holds when $B_{N}$ is replaced by $\lfloor a_{N}\rfloor$.
Proof.
Lemma 39 shows that $H_{N}\geq\frac{1}{3}H_{N+1}\geq\frac{N}{3}$. Then, by a calculation similar to (90), we obtain
(93) $\lfloor\delta^{\alpha}_{\min}\rfloor(N)\leq\frac{\log_{2}B_{N}}{H_{N-1}\log(\frac{N-1}{3})\log_{2}(\frac{N-1}{3})}\leq\frac{3\log_{2}B_{N}}{H_{N}\log(\frac{N}{4})\log_{2}(\frac{N}{4})}$
for almost all $\alpha\in[0,1]$ and for infinitely many $N$. ∎
Corollary 42.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive integers. Then, for almost all $\alpha\in[0,1]$, we have
(94) $\delta^{\alpha}_{\min}(N)\leq\frac{4\log_{2}\overline{B}_{N}}{D_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}\text{ for infinitely many $N$,}$
where $\overline{B}_{N}:=\max(a_{1},\ldots,a_{N})$, and $D_{N}=\#(A_{N}-A_{N})^{+}$.
Proof.
Since for a sequence $(a_{n})_{n\geq 1}$ of distinct positive integers we have $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq 1$, and $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\delta_{\min}^{\alpha}(N)$, we recover [ABM, Theorem 1 (Equation 5)] from Theorem 6. ∎
Remark 43.
We observe that although in [ABM, Theorem 1 (Equation 5)], the result is stated as
(95) $\delta^{\alpha}_{\min}(N)\leq\frac{\log_{2}a_{N}}{D_{N}\log N\log_{2}N}\text{ for infinitely many }N,$
we believe that for any sequence of distinct positive integers $(a_{n})_{n\geq 1}$, the correct bound should be
(96) $\delta^{\alpha}_{\min}(N)\leq\frac{4\log_{2}\max(a_{1},\ldots,a_{N})}{D_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}\text{ for infinitely many }N.$
Similarly, for any increasing sequence of distinct positive integers $(a_{n})_{n\geq 1}$, the correct bound should be
(97) $\delta^{\alpha}_{\min}(N)\leq\frac{3\log_{2}a_{N}}{D_{N}\log(\frac{N}{4})\log_{2}(\frac{N}{4})}\text{ for infinitely many }N.$
These are derived from Corollaries 42 and 41, respectively.
Remark 44.
Let $T_{N}=\#(\lfloor A_{N}-A_{N}\rfloor)^{+}$ and $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$ be defined as in Section 2.1. Since $|a_{m}-a_{n}|\geq 1$ for all positive integers $m,n$ with $m\neq n$, both sets $(\lfloor A_{N}-A_{N}\rfloor)^{+}$ and $\lfloor(A_{N}-A_{N})^{+}\rfloor$ are equal to $\{\lfloor a_{m}-a_{n}\rfloor:a_{m}>a_{n}\}$, and we obtain $T_{N}=H_{N}$.
Proposition 45.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers such that for all $N\in\mathbb{N}$. Then, for almost all $\alpha\in[0,1]$, we have
(98) $\lfloor\widetilde{\delta}_{\min}^{\alpha}\rfloor(N)\leq\frac{4\log_{2}\widetilde{B}_{N}}{H_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}\text{ for infinitely many }N,$
where $\lfloor\widetilde{\delta}_{\min}^{\alpha}\rfloor(N)=\displaystyle\min_{1\leq m\neq n\leq N}\left\|\alpha\left\lfloor a_{m}-a_{n}\right\rfloor\right\|$, $\widetilde{B}_{N}=\displaystyle\max_{1\leq m\neq n\leq N}|\left\lfloor a_{m}-a_{n}\right\rfloor|$, and $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$. Since $\widetilde{B}_{N}\leq\max(\lfloor a_{1}\rfloor,\ldots,\lfloor a_{N}\rfloor)+1=\lfloor\max(a_{1},\ldots,a_{N})\rfloor+1$, the above statement also holds when $\widetilde{B}_{N}$ is replaced by $\lfloor\max(a_{1},\ldots,a_{N})\rfloor+1$.
Proof.
Remark 44 ensures that $(\lfloor A_{N}-A_{N}\rfloor)^{+}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor(=H_{N})$. Let $(Y_{n})_{n\geq 1}$ be a sequence of distinct positive integers such that
(99) $\{Y_{n}:1\leq n\leq H_{N}\}=(\lfloor A_{N}-A_{N}\rfloor)^{+}\quad\text{for all }N.$
Mimicking the proof of Theorem 6 with a similar argument by replacing $(Z_{n},H_{N})$ with $(Y_{n},H_{N})$, we obtain
(100) $\min_{1\leq m\neq n\leq N}\left\|\alpha\left\lfloor a_{m}-a_{n}\right\rfloor\right\|=\displaystyle\min_{1\leq m\neq n\leq N}\left\|\alpha|\left\lfloor a_{m}-a_{n}\right\rfloor|\right\|=\min_{t\leq H_{N}}\|\alpha Y_{t}\|\leq\frac{4\log_{2}\widetilde{B}_{N}}{H_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}$
for infinitely many $N$. ∎
Remark 46.
For any $x\in\mathbb{R},$
[…]
If $a_{m}-a_{n}\notin\mathbb{Z}$ for all $m\neq n$, we have
(102) $\min_{1\leq m\neq n\leq N}\left\|\alpha\left(\lfloor a_{m}-a_{n}\right\rfloor+1)\right\|=\min_{1\leq m\neq n\leq N}\left\|\alpha\left(\lfloor a_{n}-a_{m}\right\rfloor+1)\right\|=$ $\min_{1\leq m\neq n\leq N}\left\|\alpha\left(-\lfloor a_{m}-a_{n}\right\rfloor)\right\|$
(103) $=$ $\displaystyle\min_{1\leq m\neq n\leq N}\left\|\alpha\left\lfloor a_{m}-a_{n}\right\rfloor\right\|$
(104) $\leq$ $\frac{4\log_{2}\widetilde{B}_{N}}{H_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}$
for infinitely many $N$.
Proposition 47.
Let $(a_{n})_{n\geq 1}$ be a sequence of distinct positive real numbers. For any integer $N\geq 2$, define
(105) $B_{N}=\max_{1\leq m\neq n\leq N}\lfloor|a_{m}-a_{n}|\rfloor\quad\text{and}\quad\widetilde{B}_{N}=\max_{1\leq m\neq n\leq N}|\lfloor a_{m}-a_{n}\rfloor|.$
Then $\widetilde{B}_{N}\geq B_{N}$ always holds.
Furthermore, the equality $\widetilde{B}_{N}=B_{N}$ is valid if and only if the maximum value $\widetilde{B}_{N}$ is achieved by a pair of indices $(m,n)$ for which the difference $|a_{m}-a_{n}|$ is an integer. In this case, $|\lfloor a_{m}-a_{n}\rfloor|=\lfloor|a_{m}-a_{n}|\rfloor$.
Proof.
Let $m,n$ be distinct indices.
  • If $a_{m}\geq a_{n}$, then
    (106) $|\lfloor a_{m}-a_{n}\rfloor|=\lfloor a_{m}-a_{n}\rfloor=\lfloor|a_{m}-a_{n}|\rfloor.$
  • If $a_{m}<a_{n}$, then using the identities $\lfloor-x\rfloor=-\lceil x\rceil$ and $\lceil x\rceil\geq\lfloor x\rfloor$ for real $x$, we obtain
    (107) $|\lfloor a_{m}-a_{n}\rfloor|=|\lfloor-(a_{n}-a_{m})\rfloor|=|-\lceil a_{n}-a_{m}\rceil|=\lceil a_{n}-a_{m}\rceil\geq\lfloor a_{n}-a_{m}\rfloor=\lfloor|a_{m}-a_{n}|\rfloor,$
    where the equality holds if and only if $a_{n}-a_{m}$ is an integer. Otherwise, it is a strict inequality.
Therefore, for each pair of distinct indices $m,n$, we have $|\lfloor a_{m}-a_{n}\rfloor|\geq\lfloor|a_{m}-a_{n}|\rfloor$, and the inequality is strict if $a_{n}-a_{m}$ is not an integer.
Taking the maximum over all pairs gives
(108) $\widetilde{B}_{N}=\max_{1\leq m\neq n\leq N}|\lfloor a_{m}-a_{n}\rfloor|\geq\max_{1\leq m\neq n\leq N}\lfloor|a_{m}-a_{n}|\rfloor=B_{N}.$
4.3. Proof of Theorem 9
Theorem 6 depends on the magnitude of $a_{N}$. To establish Theorem 9 (that is, to obtain a bound independent of $a_{N}$), we modify the coprime approach of (77) by removing only those subintervals centred at $a/Z_{n}$ such that $a$ and $Z_{n}$ share a small prime factor. This ensures that the removed measure depends solely on $N$, not $a_{N}$, while sufficiently reducing the problematic overlaps responsible for the growth of the GCD sum.
4.3.1. Overlap estimates and auxiliary lemmas
Following these heuristic arguments, we now proceed to the formal proof. Let the sequence $\left(Z_{n}\right)_{n\geq 1}$ be defined as in the previous section. For the remainder of this section, we restrict our indices $m$ and $n$ in a range $(2^{k/2},2^{k}]$ for some positive integer $k$ (where, for simplicity of writing, we assume that $k$ is even). We set
$\psi(n)=\frac{1}{4n},\quad n\geq 1,$
and
$S_{n}^{*}=[0,1]\cap\left(\bigcup_{\begin{subarray}{c}0\leq a\leq Z_{n}\\ p\mid\left(a,Z_{n}\right)\Rightarrow p>4^{k}\end{subarray}}\left(\frac{a}{Z_{n}}-\frac{\psi(n)}{Z_{n}},\frac{a}{Z_{n}}+\frac{\psi(n)}{Z_{n}}\right)\right).$
To establish Theorem 9, it suffices to prove that almost every $\alpha$ belongs to infinitely many $S_{n}^{*}$. We now state the necessary lemmas, providing proofs for those for which no reference is given.
Lemma 48.
For all $n\in(2^{k/2},2^{k}],$
$\lambda(S_{n}^{*})\geq\frac{1}{n}\frac{1}{e^{\gamma}\log\left(4^{k}\right)}(1+o(1))$
as $k\rightarrow\infty$, where $\gamma$ denotes the Euler-Mascheroni constant.
Lemma 49 (Pollington-Vaughan style overlap estimate).
Let $m,n\in(2^{k/2},2^{k}]$ with $m\neq n$. Set
(109) $D(Z_{m},Z_{n})=\frac{\max\left(Z_{m}\psi(n),Z_{n}\psi(m)\right)}{\gcd(Z_{m},Z_{n})}.$
Furthermore, when $D(Z_{m},Z_{n})\geq 1$, then set
(110) $P(Z_{m},Z_{n})=\prod_{\begin{subarray}{c}p\mid\frac{Z_{m}Z_{n}}{(Z_{m},Z_{n})^{2}}\\ D(Z_{m},Z_{n})<p\leq 4^{k}\end{subarray}}\left(1+\frac{1}{p}\right),$
where the product ranges over all primes $p$ in the specified range. When $D(Z_{m},Z_{n})<1$, then set $P(Z_{m},Z_{n})=0$. Then
(111) $\lambda(S_{m}^{*}\cap S_{n}^{*})\ll\frac{\sqrt{\psi(m)\psi(n)}}{4^{k}}+P(Z_{m},Z_{n})\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*}).$
Lemma 50 (Chung-Erdős inequality).
Let $\mathcal{A}_{m},1\leq m\leq M$, be events in a probability space $(\Omega,\mathcal{F},\mathbb{P})$, such that $\mathbb{P}\left(\mathcal{A}_{m}\right)>0$ for at least one value of $m$. Then
$\mathbb{P}\left(\bigcup_{m=1}^{M}\mathcal{A}_{m}\right)\geq\frac{\left(\displaystyle\sum_{m=1}^{M}\mathbb{P}\left(\mathcal{A}_{m}\right)\right)^{2}}{\displaystyle\sum_{1\leq m,n\leq M}\mathbb{P}\left(\mathcal{A}_{m}\cap\mathcal{A}_{n}\right)}.$
Proof.
See [chandra2012borel, Theorem 1.4.3d]. ∎
Lemma 51 (Cassels’ zero-one law [Cassels_1950]).
Let $(\xi(m))_{m\geq 1}$ be a sequence of non-negative real numbers. For $m\geq 1$, set $\mathcal{A}_{m}=\displaystyle\bigcup_{0\leq a\leq m}\left(\frac{a}{m}-\frac{\xi(m)}{m},\frac{a}{m}+\frac{\xi(m)}{m}\right)$. Let $\mathcal{A}$ be the set of those $\alpha\in[0,1]$, which are contained in infinitely many $\mathcal{A}_{m}$. Then the Lebesgue measure of $\mathcal{A}$ is either 0 or 1.
Finally, we use the following lemma, which played a crucial role in Koukoulopoulos and Maynard’s proof of the Duffin-Schaeffer conjecture.
Lemma 52.
Let $(\eta(q))_{q\geq 1}$ be a sequence of real numbers in $[0,1/2]$. Set $M(q,r)=\max(q\eta(r),r\eta(q))$. Assume that there are $X<Y$ such that
(112) $\sum_{X\leq q\leq Y}\frac{\eta(q)\varphi(q)}{q}\leq 2.$
For $t\geq 1$, set
$L_{t}(q,r)=\sum_{\begin{subarray}{c}p\mid qr/\gcd(q,r)^{2}\\ p\geq t\end{subarray}}\frac{1}{p}$
and
$\mathcal{E}_{t}=\left\{(q,r)\in(\mathbb{Z}\cap[X,Y))^{2}:\gcd(q,r)\geq t^{-1}M(q,r)\text{ and }L_{t}(q,r)\geq 10\right\}.$
Then
(113) $\sum_{(q,r)\in\mathcal{E}_{t}}\frac{\eta(q)\varphi(q)}{q}\frac{\eta(r)\varphi(r)}{r}\ll\frac{1}{t}.$
Proof.
See [KoukouMaynard, Proposition 5.4]. ∎
Remark 53.
As noted in [ABM], although the above lemma is formulated with the additional requirement that the sum in (112) is at least $1$, this requirement is unnecessary and not used in the proof of [KoukouMaynard, Proposition 5.4].
We now prove Lemmas 48 and 49.
Proof of Lemma 48. Using $\psi(n)=1/2n$, we have
$\lambda(S_{n}^{*})$ $=\frac{2\psi(n)}{Z_{n}}\#\left\{0\leq a\leq Z_{n}:p\mid\left(a,Z_{n}\right)\Rightarrow p>4^{k}\right\}$
$\geq\frac{1}{nZ_{n}}Z_{n}\prod_{p\leq 4^{k}}\left(1-\frac{1}{p}\right)$
$=\frac{1}{n}\frac{1}{e^{\gamma}\log 4^{k}}(1+o(1)),$
where the product over primes is estimated using Mertens’s third theorem [Hardy, Theorem 429].
Proof of Lemma 49. The following lemma adapts the overlap estimate of Pollington and Vaughan [pollingtonvaughan1990]. For notational convenience, we use the following notation throughout this proof
$\min]=\min\left(\frac{\psi(m)}{Z_{m}},\frac{\psi(n)}{Z_{n}}\right),\quad[\max]=\max\left(\frac{\psi(m)}{Z_{m}},\frac{\psi(n)}{Z_{n}}\right).$
The Lebesgue measure of $S_{m}^{*}\cap S_{n}^{*}$ is bounded above by (see [Harman1998, p. 39])
${[\min]}\sum_{\begin{subarray}{c}1\leq a\leq Z_{m}\\ p\mid\left(a,Z_{m}\right)\Rightarrow p>4^{k}\end{subarray}}\sum_{\begin{subarray}{c}1\leq a\leq Z_{n}\\ p\mid\left(b,Z_{n}\right)\Rightarrow p>4^{k}\end{subarray}}\mathds{1}\left(\left|\frac{a}{Z_{m}}-\frac{b}{Z_{n}}\right|\leq 2[\max]\right).$
Following Pollington and Vaughan’s method, we bound the contribution from distinct fractions $a/Z_{m}\neq b/Z_{n}$, noting that we only sifted primes below $4^{k}$. Let $u_{p}=\nu_{p}(Z_{m})$ and $v_{p}=\nu_{p}(Z_{n})$. Therefore,
$Z_{m}=\prod_{p}p^{u_{p}}\quad\text{ and }\quad Z_{n}=\prod_{p}p^{v_{p}}.$
Let
$f=\prod_{\begin{subarray}{c}p\\ u_{p}=v_{p}\end{subarray}}p^{u_{p}},\quad g=\prod_{\begin{subarray}{c}p\\ u_{p}\neq v_{p}\end{subarray}}p^{\min(u_{p},v_{p})},\quad\text{ and }\quad h=\prod_{\begin{subarray}{c}p\\ u_{p}\neq v_{p}\end{subarray}}p^{\max(u_{p},v_{p})}.$
Note that
$\frac{h}{g}=\prod_{\begin{subarray}{c}p\\ u_{p}\neq v_{p}\end{subarray}}p^{\max(u_{p},v_{p})-\min(u_{p},v_{p})}=\frac{\prod_{p}p^{\max(u_{p},v_{p})}}{\prod_{p}p^{\min(u_{p},v_{p})}}=\frac{\operatorname{lcm}(Z_{m},Z_{n})}{\gcd(Z_{m},Z_{n})}=\frac{Z_{m}Z_{n}}{\gcd(Z_{m},Z_{n})^{2}}.$
We now use the representation
$D(Z_{m},Z_{n})=[\max]fh$
to obtain
$\min]gfD(Z_{m},Z_{n})=\psi(m)\psi(n).$
Therefore, as mentioned in [ABM], an adaptation of the Pollington-Vaughan method yields
(114) $\text{ [min] }\sum_{\begin{subarray}{c}1\leq a\leq Z_{m}\\ p\mid\left(a,Z_{m}\right)\Rightarrow p>4^{k}\end{subarray}}\sum_{\begin{subarray}{c}1\leq b\leq Z_{n}\\ p\mid\left(b,Z_{n}\right)\Rightarrow p>4^{k}\end{subarray}}\mathds{1}\left(0<\left|\frac{a}{Z_{m}}-\frac{b}{Z_{n}}\right|\leq 2[\max]\right)$
(115) $\ll$ $[\min]g\prod_{\begin{subarray}{c}p\mid g\\ p\leq 4^{k}\end{subarray}}\left(1-\frac{1}{p}\right)f\prod_{\begin{subarray}{c}p\mid f\\ p\leq 4^{k}\end{subarray}}\left(1-\frac{1}{p}\right)^{2}\sum_{\begin{subarray}{c}0<j\leq 2D(Z_{m},Z_{n})\\ p\mid(j,h)\Rightarrow p>4^{k}\end{subarray}}\prod_{\begin{subarray}{c}p\mid(j,f)\\ p\leq 4^{k}\end{subarray}}\left(1+\frac{1}{p}\right)$
(116) $\ll$ $\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*})P(Z_{m},Z_{n}),$
where the sum over $j$ is estimated by Brun’s sieve, mirroring the technique employed by Pollington and Vaughan as in [pollingtonvaughan1990, p. 195–196].
In the Duffin-Schaeffer framework of Pollington and Vaughan, the complete co-primality condition ensures that there is no contribution to the overlap $S_{m}^{*}\cap S_{n}^{*}$ from pairs $(a,b)$ with $a/Z_{m}=b/Z_{n}$. In our setting, these overlaps contribute
(117) $\ll[\mathrm{min}]\sum_{\begin{subarray}{c}0\leq a\leq Z_{m}\\ p\mid\left(a,Z_{m}\right)\Rightarrow p>4^{k}\end{subarray}}\sum_{\begin{subarray}{c}0\leq b\leq Z_{n}\\ p\mid\left(b,Z_{n}\right)\Rightarrow p>4^{k}\end{subarray}}\mathds{1}\left(\frac{a}{Z_{m}}=\frac{b}{Z_{n}}\right).$
We have $a/Z_{m}=b/Z_{n}$ if and only if
$a=j\frac{Z_{m}}{\gcd(Z_{m},Z_{n})}\quad\text{ and }\quad b=j\frac{Z_{n}}{\gcd(Z_{m},Z_{n})}$
for some $j$ in $0\leq j\leq\gcd(Z_{m},Z_{n})$. We claim that for all $Z_{m}$ and $Z_{n}$ for which there exists $p\leq 4^{k}$ such that the $p$-adic valuation of $Z_{m}$ is different from that of $Z_{n}$, there is no admissible value of $j$, and the double sum above is empty. In other words, a non-vanishing contribution requires that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$. However, if $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p>4^{k}$ as well, then $Z_{m}=Z_{n}$, which is a contradiction. Therefore, a non-vanishing contribution implies that there exists a prime $p>4^{k}$ such that $\nu_{p}(Z_{m})\neq\nu_{p}(Z_{n})$.
We now prove the claim mentioned above by contrapositive. Suppose that there exists an admissible value of $j$, which means that there exist $a$ and $b$ satisfying $0\leq a\leq Z_{m}$, $p\mid(a,Z_{m})\implies p>4^{k}$, and $0\leq b\leq Z_{n}$, $p\mid(b,Z_{n})\implies p>4^{k}$ such that $\frac{a}{Z_{m}}=\frac{b}{Z_{n}}$, that is, $aZ_{n}=bZ_{m}$. This shows that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$. If not, there exists $q\leq 4^{k}$ such that $\nu_{q}(Z_{m})\neq\nu_{q}(Z_{n})$. Without loss of generality, assume $\nu_{q}(Z_{n})>\nu_{q}(Z_{m})\geq 0$, that is, $\nu_{q}(Z_{n})-\nu_{q}(Z_{m})\geq 1$. Then $q$ divides $aq^{\nu_{q}(Z_{n})-\nu_{q}(Z_{m})}\displaystyle\prod_{p\neq q}p^{\nu_{p}(Z_{n})}=b\displaystyle\prod_{p\neq q}p^{\nu_{p}(Z_{m})}$, that is, $q$ divides $b$. In addition, $\nu_{q}(Z_{n})\geq 1$ implies $q|Z_{n}$, so $q|(b,Z_{n})$ with $q\leq 4^{k}$, which is a contradiction. This contradiction shows that $\nu_{p}(Z_{m})=\nu_{p}(Z_{n})$ for all $p\leq 4^{k}$.
Since a non-vanishing contribution implies the existence of a prime $p>4^{k}$ such that $\nu_{p}(Z_{m})\neq\nu_{p}(Z_{n})$, we have $\gcd(Z_{m},Z_{n})\leq\min(Z_{m},Z_{n})/4^{k}$. Thus, we can estimate (117) by
$\ll\min\left(\frac{\psi(m)}{Z_{m}},\frac{\psi(n)}{Z_{n}}\right)\frac{\min(Z_{m},Z_{n})}{4^{k}}\ll\frac{\sqrt{\psi(m)\psi(n)}}{\sqrt{Z_{m}Z_{n}}}\frac{\sqrt{Z_{m}Z_{n}}}{4^{k}}\ll\frac{\sqrt{\psi(m)\psi(n)}}{4^{k}}.$
This together with (114) proves the lemma.
4.3.2. Proof of Theorem 9, upper bound independent of the size of $a_{N}$.
Assume that $k$ is a fixed, even, and “large” number. We decompose all numbers $Z_{n}$ as follows.
$Z_{n}=Z_{n}^{\text{smooth }}\cdot Z_{n}^{\text{rough }},\quad 2^{k/2}<n\leq 2^{k},$
where $Z_{n}^{\text{smooth }}$ has only prime factors of size at most $4^{k}$, and $Z_{n}^{\text{rough }}$ has only prime factors of size larger than $4^{k}$, that is, $Z_{n}^{\text{smooth }}=\displaystyle\prod_{p\leq 4^{k}}p^{\nu_{p}(Z_{n})}$ and $Z_{n}^{\text{rough }}=\displaystyle\prod_{p>4^{k}}p^{\nu_{p}(Z_{n})}$. Write
$\left\{b_{1},\ldots,b_{H}\right\}=\left\{Z_{n}^{\text{rough }},2^{k/2}<n\leq 2^{k}\right\}$
for some appropriate $H$, where we assume that $b_{1},\ldots,b_{H}$ are sorted in increasing order. Clearly, $H\leq 2^{k}-2^{k/2}$, but $H$ might actually be smaller since we could have $Z_{m}^{\text{rough }}=Z_{n}^{\text{rough }}$ for some $m\neq n$. For every $n\in(2^{k/2},2^{k}]$, we now define
$Y_{n}\coloneqq Z_{n}^{\text{smooth }}p_{h}^{(k)},$
where $h$ is the uniquely defined index for which $Z_{n}^{\text{rough }}=b_{h}$, and where $p_{1}^{(k)}$ denotes the smallest prime exceeding $4^{k},p_{2}^{(k)}$ denotes the second-smallest prime exceeding $4^{k}$, and so on. The point in the construction of the numbers $\left(Y_{n}\right)_{2^{k/2}<n\leq 2^{k}}$ is that on the one hand, in Lemma 49, we can replace $P(Z_{m},Z_{n})$ and $D(Z_{m},Z_{n})$ by $P\left(Y_{m},Y_{n}\right)$ and $D\left(Y_{m},Y_{n}\right)$ in the relevant situations, since the small prime factors of $Z_{m}$ and $Z_{n}$ are the same as those of $Y_{m}$ and $Y_{n}$, respectively. On the other hand, we have
$\prod_{p\mid Y_{n}}\left(1-\frac{1}{p}\right)\leq\prod_{\begin{subarray}{c}p\mid Z_{n}\\ p\leq 4^{k}\end{subarray}}\left(1-\frac{1}{p}\right)\leq\left(1-\frac{1}{4^{k}}\right)^{-1}\prod_{p\mid Y_{n}}\left(1-\frac{1}{p}\right),$
where we used that by construction the “small” prime factors of $Z_{n}$ and $Y_{n}$ coincide, and $Y_{n}$ has an additional prime factor that exceeds $4^{k}$. Consequently,
(118) $\frac{2\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\leq\lambda(S_{n}^{*})\leq\frac{2\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\left(1-\frac{1}{4^{k}}\right)^{-1}.$
Thus, we can control the size of the Euler totient function of $Y_{n}$ (while we cannot control it for $Z_{n}$, which might have many large prime factors). This will allow us to apply Lemma 52. By Lemma 50, we have
(119) $\lambda\Big(\bigcup_{2^{k/2}<n\leq 2^{k}}S_{n}^{*}\Big)\geq\frac{\Big(\displaystyle\sum_{2^{k/2}<n\leq 2^{k}}\lambda(S_{n}^{*})\Big)^{2}}{\displaystyle\sum_{2^{k/2}<m,n\leq 2^{k}}\lambda(S_{m}^{*}\cap S_{n}^{*})}.$
By Lemma 48, we have
(120) $\sum_{2^{k/2}<n\leq 2^{k}}\lambda(S_{n}^{*})\geq\sum_{2^{k/2}<n\leq 2^{k}}\frac{1}{n}\frac{1}{e^{\gamma}\log 4^{k}}(1+o(1))\geq 0.14$
for sufficiently large $k$. On the other hand, we can assume without loss of generality that
(121) $\sum_{2^{k/2}<n\leq 2^{k}}\lambda(S_{n}^{*})\leq 0.99$
(if the sum of measures is even larger, we can just delete some of the sets $S_{n}^{*}$ ). Thus, we can control the size of the numerator on the right-hand side of (119). To estimate the denominator of the right-hand side of (119), by Lemma 49, we have
(122) $\lambda(S_{m}^{*}\cap S_{n}^{*})\ll\frac{\sqrt{\psi(m)\psi(n)}}{4^{k}}+P(Z_{m},Z_{n})\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*}).$
Trivially,
(123) $\sum_{2^{k/2}<m,n\leq 2^{k}}\frac{\sqrt{\psi(m)\psi(n)}}{4^{k}}\ll 1.\$
Note that whenever $Z_{m}^{\text{rough }}\neq Z_{n}^{\text{rough }}$, we have $\gcd(Z_{m},Z_{n})\leq 4^{-k}\min(Z_{m},Z_{n})$, and so
$D(Z_{m},Z_{n})=\frac{\max\left(Z_{m}\psi(n),Z_{n}\psi(m)\right)}{\gcd(Z_{m},Z_{n})}\geq\frac{4^{k}\max(Z_{m},Z_{n})}{2^{k}\min(Z_{m},Z_{n})}\geq 2^{k}.$
Therefore, in this case
$P(Z_{m},Z_{n})\ll\prod_{2^{k}\leq p\leq 4^{k}}\left(1+\frac{1}{p}\right)\ll 1.$
On the other hand, if $Z_{m}^{\text{rough }}=Z_{n}^{\text{rough }}$, then it is easily seen that $D\left(Y_{m},Y_{n}\right)=D(Z_{m},Z_{n})$. In the next displayed formula, all sums are taken over $m,n$ in the range $2^{k/2}<m,n\leq 2^{k}$. Using (118), we can estimate
(124) $\sum_{m,n}P(Z_{m},Z_{n})\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*})$
(125) $\ll$ $\sum_{\begin{subarray}{c}m,n\\ Z_{m}^{\text{rough }}\neq Z_{n}^{\text{rough }}\end{subarray}}\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*})+\sum_{\begin{subarray}{c}m,n\\ Z_{m}^{\text{rough }}=Z_{n}^{\text{rough }}\end{subarray}}\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*})\prod_{\begin{subarray}{c}p\mid\frac{Z_{m}Z_{n}}{(Z_{m},Z_{n})^{2}}\\ D(Z_{m},Z_{n})<p\leq 4^{k}\end{subarray}}\left(1+\frac{1}{p}\right)$
(126) $\ll$ $\sum_{m,n}\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*})+\sum_{m,n}\frac{\psi(m)\varphi\left(Y_{m}\right)}{Y_{m}}\frac{\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\prod_{\begin{subarray}{c}p\mid\frac{Y_{m}Y_{n}}{\left(Y_{m},Y_{n}\right)^{2}}\\ D\left(Y_{m},Y_{n}\right)<p\leq 4^{k}\end{subarray}}\left(1+\frac{1}{p}\right)$
(127) $\ll$ $\sum_{m,n}\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*})+\sum_{t=0}^{k-1}\sum_{\begin{subarray}{c}2^{k/2}<m,n\leq 2^{k}\\ D\left(Y_{m},Y_{n}\right)\in[4^{t},4^{t+1})\end{subarray}}\frac{\psi(m)\varphi\left(Y_{m}\right)}{Y_{m}}\frac{\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\prod_{\begin{subarray}{c}p\mid\frac{Y_{m}Y_{n}}{\left(Y_{m},Y_{n}\right)^{2}}\\ p>4^{t+1}\end{subarray}}\left(1+\frac{1}{p}\right).$
We divide the second sum in the last line of (124) depending on whether $\left(Y_{m},Y_{n}\right)$ is in $\mathcal{E}_{4^{t+1}}$ or not. Bounding the Euler product trivially, we obtain
(128) $\sum_{t=0}^{k-1}\sum_{\begin{subarray}{c}2^{k/2}<m,n\leq 2^{k},\\ D\left(Y_{m},Y_{n}\right)\in[4^{t},4^{t+1}),\\ \left(Y_{m},Y_{n}\right)\notin\mathcal{E}_{{4}^{t+1}}\end{subarray}}\frac{\psi(m)\varphi\left(Y_{m}\right)}{Y_{m}}\frac{\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\prod_{\begin{subarray}{c}p\mid\frac{Y_{m}Y_{n}}{\left(Y_{m},Y_{n}\right)^{2}}\\ p>4^{t+1}\end{subarray}}\left(1+\frac{1}{p}\right)$
(129) $\ll$ $\sum_{t=0}^{k-1}\sum_{\begin{subarray}{c}2^{k/2}<m,n\leq 2^{k},\\ D\left(Y_{m},Y_{n}\right)\in[4^{t},4^{t+1})\end{subarray}}\frac{\psi(m)\varphi\left(Y_{m}\right)}{Y_{m}}\frac{\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}$
(130) $\ll$ $\sum_{m,n}\lambda\left(S_{m}^{*}\right)\lambda(S_{n}^{*}).$
In the case where $\left(Y_{m},Y_{n}\right)\in\mathcal{E}_{4^{t+1}}$, we apply Lemma 52 with
(131) $\eta(q)=\begin{cases}\psi(n)&\quad\text{if }q=Y_{n}\text{ for some }n\in(2^{k/2},2^{k}],\\ 0&\quad\text{otherwise,}\end{cases}$
and with $X$ and $Y$ defined as the minimum and maximum, respectively, of the set $\{Y_{n}:2^{k/2}<n\leq 2^{k}\}$. Then by (120), and (121), we have
(132) $\sum_{X\leq q\leq Y}\frac{\eta(q)\varphi(q)}{q}=\sum_{2^{k/2}<n\leq 2^{k}}\frac{\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\in[1/16,1/2],$
for sufficiently large $k$.
In the same way as Koukoulopoulos and Maynard deduce [KoukouMaynard, Theorem 1] from Lemma 52, we obtain
(133) $\sum_{t=0}^{k-1}\sum_{\begin{subarray}{c}2^{k/2}<m,n\leq\leq 2^{k},\\ D\left(Y_{m},Y_{n}\right)\leq 4^{t+1},\\ (Y_{m},Y_{n})\in\mathcal{E}_{4}t+1\end{subarray}}\frac{\psi(m)\varphi\left(Y_{m}\right)}{Y_{m}}\frac{\psi(n)\varphi\left(Y_{n}\right)}{Y_{n}}\prod_{\begin{subarray}{c}p\mid\frac{Y_{m}Y_{n}}{\left(Y_{m},Y_{n}\right)^{2}}\\ p>4^{t+1}\end{subarray}}\left(1+\frac{1}{p}\right)\ll\sum_{t=0}^{k-1}\frac{t}{4^{t}}\ll 1.$
Combining (122), (123), and (124), we obtain
$\sum_{2^{k/2}\leq m,n\leq 2^{k}}\lambda(S_{m}^{*}\cap S_{n}^{*})\ll 1.$
Thus, by (119) and (120), for all sufficiently large $k$,
$\lambda\left(\bigcup_{2^{k/2}<n\leq 2^{k}}S_{n}^{*}\right)\gg 1$
where the implied constant is independent of $k$. Since $k$ can be chosen arbitrarily large, this implies
$\lambda\left(\bigcap_{\ell=1}^{\infty}\left(\bigcup_{n=\ell}^{\infty}S_{n}^{*}\right)\right)>0,$
and since $S_{n}^{*}\subset S_{n}$, we clearly also have
$\lambda\left(\bigcap_{\ell=1}^{\infty}\left(\bigcup_{n=\ell}^{\infty}S_{n}\right)\right)>0.$
Thus, the measure of the limsup set is positive, which, by Cassels’ zero-one law (Lemma 51), implies that it is indeed $1$. In other words, almost all $\alpha\in[0,1]$ are contained in infinitely many sets $S_{n}$. Thus, for almost all $\alpha$, there are infinitely many $n$ such that
$\left\|\alpha Z_{n}\right\|\leq\frac{1}{4n}.$
By the construction of $Z_{n}$, for every $n$, there is a unique $N$ such that $H_{N-1}<n\leq H_{N}$, and Lemma 37 shows that $n>H_{N-1}\geq\frac{1}{4}H_{N}$. Therefore, for all $n\in\mathbb{N}$ with $n\in(H_{N-1},H_{N}]$, we have
$\left\|\alpha Z_{n}\right\|\leq\frac{1}{4n}\leq\frac{1}{H_{N}},$
which implies that
(134) $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=$ $\displaystyle\min_{1\leq m\neq n\leq N}\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|=\displaystyle\min_{t\leq H_{N}}\|\alpha Z_{t}\|\leq\|\alpha Z_{n}\|\leq\frac{1}{H_{N}}.$
Thus,
(135) $\lfloor\delta^{\alpha}_{\min}\rfloor(N)\leq\frac{1}{H_{N}}$
for infinitely many $N$ and for almost all $\alpha$, as required.
4.4. Proof of Theorem 11
To prove Theorem 11, we use the methodology developed by Rudnick [Rudnick]. Let $A_{N}$ be defined as in Section 2.1. Denote the set of floored positive differences by $\mathcal{Z}_{N}$, that is, $\mathcal{Z}_{N}=\lfloor(A_{N}-A_{N})^{+}\rfloor=\{Z_{1},\dots,Z_{H_{N}}\}$. Therefore, $|\mathcal{Z}_{N}|=H_{N}$ and the hypothesis ensure that $\mathcal{Z}_{N}\subseteq\mathbb{N}$.
Let $f\in C_{c}^{\infty}(\mathbb{R})$ be a non-negative, smooth, even function supported in [1/2,1/2][-1/2,1/2] with $\int_{-\infty}^{\infty}f(x)dx=1$. Define the function
(136) $F_{M}(x)=\sum_{j\in\mathbb{Z}}f(M(x+j)),$
which is $1$-periodic and is localised on the scale $1/M$.
We then define the counting function
(137) $D(N,M)(\alpha)=\sum_{z\in\mathcal{Z}_{N}}F_{M}(\alpha z).$
Since
(138) $\int_{0}^{1}F_{M}(\alpha z)\,d\alpha=\sum_{j\in\mathbb{Z}}\int_{0}^{1}f(M(\alpha z+j))\,d\alpha=\frac{1}{z}\sum_{j\in\mathbb{Z}}\int_{j}^{j+z}f(My)\,dy=\int_{-\infty}^{\infty}f(My)\,dy=\frac{1}{M},$
the expected value of $D(N,M)$ is
(139) $\mathbb{E}(D(N,M))=\int_{0}^{1}D(N,M)(\alpha)\,d\alpha=\sum_{z\in\mathcal{Z}_{N}}\int_{0}^{1}F_{M}(\alpha z)\,d\alpha=\frac{|\mathcal{Z}_{N}|}{M}=\frac{H_{N}}{M}.$
The Fourier expansion of $F_{M}$ is given by
(140) $F_{M}(x)=\sum_{k\in\mathbb{Z}}\frac{1}{M}\widehat{f}\left(\frac{k}{M}\right)e(kx),$
where $\widehat{f}(y)=\int_{-\infty}^{\infty}f(x)e^{-2\pi ixy}\,dx$ and $e(x)=e^{2\pi ix}$. Putting this into the definition of $D(N,M)$ yields
(141) $D(N,M)(\alpha)=\sum_{z\in\mathcal{Z}_{N}}F_{M}(\alpha z)=\sum_{k\in\mathbb{Z}}\frac{1}{M}\widehat{f}\left(\frac{k}{M}\right)\sum_{z\in\mathcal{Z}_{N}}e(k\alpha z).$
Since $\frac{\widehat{f}(0)}{M}|\mathcal{Z}_{N}|=\frac{H_{N}}{M}$, The expected value comes from the term corresponding to $k=0$. Therefore, the variance, denoted by $\text{Var }D(N,M)$, is the second moment of the sum over non-zero frequencies:
(142) $\text{Var }(D(N,M))=$ $\int_{0}^{1}|D(N,M)(\alpha)-\mathbb{E}[D(N,M)]|^{2}d\alpha$
(143) $=$ $\int_{0}^{1}\Big|\sum_{0\neq k\in\mathbb{Z}}\frac{1}{M}\widehat{f}\left(\frac{k}{M}\right)\sum_{z\in\mathcal{Z}_{N}}e(k\alpha z)\Big|^{2}d\alpha$
(144) $=$ $\sum_{k_{1},k_{2}\neq 0}\frac{1}{M^{2}}\widehat{f}\left(\frac{k_{1}}{M}\right)\widehat{f}\left(\frac{k_{2}}{M}\right)\sum_{v_{1},v_{2}\in\mathcal{Z}_{N}}\int_{0}^{1}e(\alpha(k_{1}v_{1}-k_{2}v_{2}))d\alpha.$
The integral equals $1$ if $k_{1}v_{1}-k_{2}v_{2}=0$ and 0 otherwise, so
$\int_{0}^{1}e(\alpha(k_{1}v_{1}-k_{2}v_{2}))d\alpha=\mathds{1}_{\{0\}}(k_{1}v_{1}-k_{2}v_{2}).$
As $(Z_{n})_{n\geq 1}$ consists of distinct integers, we apply the GCD sum estimate from [Bondarenko-Seip] to obtain
(145) $\sum_{m,n=1}^{H_{N}}\frac{\gcd(Z_{m},Z_{n})}{\sqrt{Z_{m}Z_{n}}}\ll H_{N}\exp\left(A\sqrt{\frac{\log H_{N}\log_{3}H_{N}}{\log_{2}H_{N}}}\right)$
for some absolute constant $A$ less than $7$. Combined with [Rudnick, Lemma 4], this yields
(146) $\text{Var }(D(N,M))=$ $\sum_{v_{1},v_{2}\in\mathcal{Z}_{N}}\sum_{k_{1},k_{2}\neq 0}\frac{1}{M^{2}}\widehat{f}\left(\frac{k_{1}}{M}\right)\widehat{f}\left(\frac{k_{2}}{M}\right)\mathds{1}_{\{0\}}(k_{1}v_{1}-k_{2}v_{2})$
(147) $\ll$ $\frac{1}{M}\sum_{v_{1},v_{2}\in\mathcal{Z}_{N}}\frac{\gcd(v_{1},v_{2})}{\sqrt{v_{1}v_{2}}}$
(148) $\ll$ $\frac{1}{M}\sum_{1\leq k,l\leq H_{N}}\frac{\gcd(Z_{k},Z_{l})}{\sqrt{Z_{k}Z_{l}}}$
(149) $\ll$ $\frac{H_{N}}{M}\exp\left(A\sqrt{\frac{\log H_{N}\log_{3}H_{N}}{\log_{2}H_{N}}}\right)$
(150) $\ll$ $\frac{H_{N}}{M}\exp\left(A\sqrt{\frac{\log N\log_{3}N}{\log_{2}N}}\right)\ll\frac{H_{N}}{M}\exp\left(\epsilon\log N\right)=\frac{H_{N}N^{\epsilon}}{M}$
for any $\epsilon>0$.
Proof of Theorem 11 (upper bound for all except finitely many $N$).
The proof is carried out in three steps. First, we establish a relationship between $H_{N_{k+1}}$ and $H_{N_{k}}$ along some subsequence $N_{k}$. In the next step, we show that the statistic $D(N,M)$ concentrates around its mean for an appropriate choice of $M$. Then, we use this concentration to deduce the existence of small gaps.
Lemma 54.
Let $\eta\in(0,2]$ be a fixed parameter and $N_{k}=\lfloor k^{\eta/2}\rfloor$. Then there exists a constant $C>0$ such that for all $k\geq 1$,
(151) $H_{N_{k+1}}\leq C_{\eta}H_{N_{k}},$
where $C_{\eta}=5+2\eta$. In particular, one can take $C_{\eta}=9$.
Proof.
With $A_{N}$ defined in Section 2.1, we have $A_{N_{k+1}}=A_{N_{k}}\cup B$, where $B$ contains $N_{k+1}-N_{k}$ new elements.
Each new element $x\in B$ creates at most $2N_{k}$ new positive differences:
  • $\lfloor x-y\rfloor$ for $y\in A_{N_{k}}$ (positive if $x>y$).
  • $\lfloor y-x\rfloor$ for $y\in A_{N_{k}}$ (positive if $y>x$).
Thus,
(152) $H_{N_{k+1}}\leq H_{N_{k}}+2N_{k}(N_{k+1}-N_{k}).$
Since $N_{k}=\lfloor k^{\eta/2}\rfloor$, we have
(153) $N_{k+1}-N_{k}\leq(k+1)^{\eta/2}-(k^{\eta/2}-1)\leq 1+[(k+1)^{\eta/2}-k^{\eta/2}].$
We now apply the mean value theorem to $g(x)=x^{\eta/2}$ on $[k,k+1]$ to obtain
(154) $(k+1)^{\eta/2}-k^{\eta/2}=\frac{\eta}{2}c^{\eta/2-1}\quad\text{for some }c\in(k,k+1).$
Since $c>k$ and $\eta\leq 2$ (so $\eta/2-1\leq 0$), we have $c^{\eta/2-1}\leq k^{\eta/2-1}$. Therefore,
(155) $(k+1)^{\eta/2}-k^{\eta/2}\leq\frac{\eta}{2}k^{\eta/2-1},$
and thus
(156) $N_{k+1}-N_{k}\leq 1+\frac{\eta}{2}k^{\eta/2-1}.$
From (152) and the lower bound $H_{N_{k}}\geq N_{k}/2$, we get
(157) $\frac{H_{N_{k+1}}}{H_{N_{k}}}\leq 1+\frac{2N_{k}(N_{k+1}-N_{k})}{H_{N_{k}}}\leq 1+\frac{2N_{k}(N_{k+1}-N_{k})}{N_{k}/2}=1+4(N_{k+1}-N_{k})\leq 5+2\eta\,k^{\eta/2-1}.$
Since $\eta\leq 2$, we have $k^{\eta/2-1}\leq 1$ for all $k\geq 1$. Therefore,
(158) $\frac{H_{N_{k+1}}}{H_{N_{k}}}\leq 5+2\eta.$
Lemma 55.
Let $\eta\in(0,2]$, and define $N_{k}=\lfloor k^{2/\eta}\rfloor$ and $M_{k}=H_{N_{k}}/N_{k}^{\eta}$, where $N_{k}$ is defined as in Lemma 54. Then, for almost every $\alpha\in[0,1]$,
(159) $D(N_{k},M_{k})(\alpha)\sim\frac{H_{N_{k}}}{M_{k}}.$
Proof.
Let $\delta>0$ be arbitrary. By Chebyshev’s inequality and the variance bound (146), we have
(160) $\mathbb{P}\left(\left|D(N_{k},M_{k})(\alpha)-\frac{H_{N_{k}}}{M_{k}}\right|>\delta\cdot\frac{H_{N_{k}}}{M_{k}}\right)$ $\leq\frac{\text{Var }D(N_{k},M_{k})}{\delta^{2}\left(\frac{H_{N_{k}}}{M_{k}}\right)^{2}}\ll\frac{\frac{H_{N_{k}}}{M_{k}}N_{k}^{\epsilon}}{\delta^{2}\frac{H_{N_{k}}^{2}}{M_{k}^{2}}}=\frac{M_{k}N_{k}^{\epsilon}}{\delta^{2}H_{N_{k}}}$
for any $\epsilon>0$. Substituting $M_{k}=H_{N_{k}}/N_{k}^{\eta}$ gives
(161) $\mathbb{P}\left(\left|\frac{D(N_{k},M_{k})}{H_{N_{k}}/M_{k}}-1\right|>\delta\right)\ll\frac{1}{\delta^{2}N_{k}^{\eta-\epsilon}}.$
Choose $0<\epsilon<\frac{\eta}{2}$, so that $\eta-\epsilon>\frac{\eta}{2}>0$. Since $\lfloor x\rfloor\geq\frac{x}{2}$ for $x>2$, we have $N_{k}\geq\frac{k^{2/\eta}}{2}$ for large $k$. With the subsequence $N_{k}=\lfloor k^{2/\eta}\rfloor$, we then have
(162) $\sum_{k=1}^{\infty}\frac{1}{N_{k}^{\eta-\epsilon}}<\infty.$
By the Borel–Cantelli lemma, for almost all $\alpha$, there exists $k_{0}(\alpha)$ such that
(163) $\left|\frac{D(N_{k},M_{k})(\alpha)}{H_{N_{k}}/M_{k}}-1\right|\leq\delta\quad\text{for all }k\geq k_{0}(\alpha),$
which establishes the corollary. ∎
Lemma 55 implies that for almost all $\alpha$, there exists $k_{0}(\alpha)$ such that for all $k\geq k_{0}(\alpha)$,
(164) $D(N_{k},M_{k})(\alpha)\sim\frac{H_{N_{k}}}{M_{k}}=N_{k}^{\eta}.$
In particular, since $N_{k}^{\eta}\to\infty$ as $k\to\infty$, we have $D(N_{k},M_{k})(\alpha)>1$ for all sufficiently large $k$.
We have shown that for almost all $\alpha$, there exists $k_{0}(\alpha)$ such that $D(N_{k},M_{k})(\alpha)>1$ for all $k\geq k_{0}(\alpha)$. It then follows from the definition of $D(N,M)(\alpha)$ (see (137)) that for each $k\geq k_{0}(\alpha)$, there exists some $z\in\mathcal{Z}_{N_{k}}$ such that $F_{M_{k}}(\alpha z)>0$, that is, there exists $j_{0}\in\mathbb{Z}$ such that $f(M_{k}(\alpha z+j_{0}))>0$.
Since $f$ is supported in [1/2,1/2][-1/2,1/2], we have $|\alpha z+j_{0}|\leq\frac{1}{2M_{k}}$ for some integer $j_{0}$, which implies $\|\alpha z\|\leq\frac{1}{M_{k}}$ for some $z\in\mathcal{Z}_{N_{k}}$ and all $k\geq k_{0}(\alpha)$. Therefore, for almost every $\alpha$, there exists $k_{0}(\alpha)$ such that
(165) $\lfloor\delta_{\min}^{\alpha}\rfloor(N_{k})=\displaystyle\min_{\widetilde{z}\in\mathcal{Z}_{N_{k}}}\|\alpha\widetilde{z}\|\leq\|\alpha z\|\leq\frac{1}{M_{k}}=\frac{N_{k}^{\epsilon}}{H_{N_{k}}}$
for all $k\geq k_{0}(\alpha)$ and any $\epsilon>0$.
Let $N\geq\lfloor k_{0}(\alpha)^{2/\eta}\rfloor$. Then $N\in\big[\lfloor(k_{0}(\alpha)+m)^{2/\eta}\rfloor,\lfloor(k_{0}(\alpha)+m+1)^{2/\eta}\rfloor\big)$ for some integer $m\geq 0$, that is, $N\in[N_{k_{0}(\alpha)+m},N_{k_{0}(\alpha)+m+1})$. Since $H_{N}$ increases in $N$, we have $H_{N_{k_{0}(\alpha)+m}}\leq H_{N}\leq H_{N_{k_{0}(\alpha)+m+1}}$. Applying Lemma 54, we obtain for all $N\geq\lfloor k_{0}(\alpha)^{2/\eta}\rfloor$,
(166) $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\lfloor\delta_{\min}^{\alpha}\rfloor(N_{k_{0}(\alpha)+m})\leq\frac{N_{k_{0}(\alpha)+m}^{\epsilon}}{H_{N_{k_{0}(\alpha)+m}}}\leq\frac{9N^{\epsilon}}{H_{N_{k_{0}(\alpha)+m+1}}}\leq\frac{9N^{\epsilon}}{H_{N}}$
for any $\epsilon>0$.
Let $b>0$. Then, for all $N\geq\max(\lfloor k_{0}(\alpha)^{2/\eta}\rfloor,9^{2/b}),$
(167) $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\frac{9N^{b/2}}{H_{N}}\leq\frac{N^{b}}{H_{N}}$
for any $b>0$.
5. Proof of Theorem 13
We now provide a proof for Theorem 13. The following statement, widely known as Catlin’s conjecture [CATLIN1976], was recently resolved as a consequence of Koukoulopoulos and Maynard’s proof of the Duffin–Schaeffer conjecture [KoukouMaynard, Theorem 2].
Lemma 56.
Let $(\psi(k))_{k\geq 1}$ be a sequence of non-negative reals. Let $\mathcal{B}$ denote the set of those $\alpha\in[0,1]$ for which the inequality
$\left|\alpha-\frac{a}{k}\right|\leq\frac{\psi(k)}{k}$
has infinitely many solutions $(a,k)$ with $0\leq a\leq k$. Then $\lambda(\mathcal{B})=0$ or $\lambda(\mathcal{B})=1$, according to whether the series
(168) $\sum_{k=1}^{\infty}\varphi(k)\sup_{b\geq 1}\left\{\frac{\psi(bk)}{bk}\right\}$
is convergent or divergent, respectively.
Lemma 57.
Let $(\lambda_{n})_{n\geq 1}$ be an arbitrary sequence of real numbers. Suppose that as a function of the index, $\lambda$ satisfies the following condition: There exist two positive constants $\epsilon$ and $c$ such that whenever the index increases from $n$ by more than $\frac{n}{(\log n)^{1+\epsilon}}$, the value of $\lambda$ increases by at least $c$. Then, the sequence $(\alpha\lambda_{n})_{n\geq 1}$ is uniformly distributed for almost all $\alpha\in[0,1)$.
Proof.
See [Weyl1916, Satz 21]. ∎
Lemma 58.
Let $(\lambda_{n})_{n\geq 1}$ be a sequence of distinct integers. Then, the sequence $(\alpha\lambda_{n})_{n\geq 1}$ is uniformly distributed mod $1$ for almost all real numbers $\alpha$.
Proof.
See [kuipers1974uniform, Theorem 4.1]. ∎
Lemma 59.
Let $(a_{n})_{n\geq 1}$ be a sequence of non-negative real numbers and $A$ be an infinite subset of $\mathbb{N}$ such that for each $k\in A$, $\sum_{\begin{subarray}{c}d|k\\ d\geq\log{k}\end{subarray}}a_{d}\geq\epsilon$ for some fixed $\epsilon>0$. Then $\sum_{k=1}^{\infty}a_{k}=\infty$.
Proof.
For each positive integer $N$, let $T(N)\coloneqq\#\{d:d\mid N,\,d\geq\log N\}$, and we enumerate the divisors as
$\{d:d\mid N,\,d\geq\log N\}=\{d_{1}^{N},d_{2}^{N},\dots,d_{T(N)}^{N}\},$
where $d_{1}^{N}<d_{2}^{N}<\ldots<d_{T(N)}^{N}$. Since $A$ is an infinite set, there exists $k_{1}\in A$ such that $\sum_{\begin{subarray}{c}d\mid k_{1}\\ d\geq\log{k_{1}}\end{subarray}}a_{d}\geq\epsilon$, that is,
(169) $a_{d_{1}^{k_{1}}}+a_{d_{2}^{k_{1}}}+\ldots+a_{d_{T(k_{1})}^{k_{1}}}\geq\epsilon.$
Now, choose $k_{2}\in A$ such that $k_{2}\geq e^{k_{1}}$ (this is possible since $A$ is unbounded). By hypothesis, we also have $\sum_{\begin{subarray}{c}d\mid k_{2}\\ d\geq\log{k_{2}}\end{subarray}}a_{d}\geq\epsilon$, that is,
(170) $a_{d_{1}^{k_{2}}}+a_{d_{2}^{k_{2}}}+\ldots+a_{d_{T(k_{2})}^{k_{2}}}\geq\epsilon.$
Now, for each $d\mid k_{2}$ satisfying $d\geq\log k_{2}$, we have $d\geq k_{1}$, that is, $d>d_{i}^{k_{1}}$ for each $i=1,\ldots,T(k_{1})$. Thus, $d_{j}^{k_{2}}>d_{i}^{k_{1}}$ for all $i\in\{1,\ldots,T(k_{1})\}$ and $j\in\{1,\ldots,T(k_{2})\}$.
By induction, construct a sequence $(k_{n})_{n\geq 1}$ in $A$ such that $k_{n}\geq e^{k_{n-1}}$ and $\sum_{\begin{subarray}{c}d\mid k_{n}\\ d\geq\log{k_{n}}\end{subarray}}a_{d}\geq\epsilon$, that is,
(171) $a_{d_{1}^{k_{n}}}+a_{d_{2}^{k_{n}}}+\ldots+a_{d_{T(k_{n})}^{k_{n}}}\geq\epsilon.$
By construction, we have $d_{i}^{k_{m}}<d_{j}^{k_{n}}$ for all $m<n$, $i\in\{1,\ldots,T(k_{m})\}$, and $j\in\{1,\ldots,T(k_{n})\}$. Also, for fixed $N$, $d_{r}^{N}<d_{s}^{N}$ for $1\leq r<s\leq T(N)$. Therefore,
(172) $\displaystyle\sum_{k=1}^{d_{T(k_{n})}^{k_{n}}}a_{k}\geq(a_{d_{1}^{k_{1}}}+\ldots+a_{d_{T(k_{1})}^{k_{1}}})+(a_{d_{1}^{k_{2}}}+\ldots+a_{d_{T(k_{2})}^{k_{2}}})+\ldots+(a_{d_{1}^{k_{n}}}+\ldots+a_{d_{T(k_{n})}^{k_{n}}})\geq n\epsilon$
for each $n\in\mathbb{N}$. Hence, $\sum_{k=1}^{\infty}a_{k}=\infty$. ∎
Corollary 60.
Let $(a_{n})_{n\geq 1}$ be a sequence of non-negative real numbers and $A$ be an infinite subset of $\mathbb{N}$ such that for all sufficiently large $k\in A$, $\sum_{\begin{subarray}{c}d|k\\ d\geq\log{k}\end{subarray}}a_{d}\geq\epsilon$ for some fixed $\epsilon>0$. Then $\sum_{k=1}^{\infty}a_{k}=\infty$.
Proof.
The result follows from Lemma 59 with a minor adjustment to the construction. Since the hypothesis applies to all sufficiently large $k\in A$, we choose $k_{1}$ large enough to satisfy the condition. For $n\geq 2$, we select $k_{n}\in A$ such that $k_{n}\geq e^{k_{n-1}}$. This ensures that all $k_{n}$ are sufficiently large, as $k_{n}>k_{n-1}$ for all $n$. The remainder of the proof proceeds identically. ∎
Lemma 61.
Let $k\in\mathbb{N}$ and let $\varphi$ be Euler’s totient function.
  1. (1)
    $\sum_{d\mid k}\varphi(d)=k$.
  2. (2)
    For all sufficiently large $k$, that is, for $k\gg 1$, $\sum_{\begin{subarray}{c}d\mid k\\ d<\log k\end{subarray}}\varphi(d)<\frac{k}{2}$.
Proof.
We only prove $(2)$, as $(1)$ is a standard fact in number theory. Since $\varphi(d)\leq d$ for all $d\in\mathbb{N}$, we have
(173) $\sum_{\begin{subarray}{c}d\mid k\\ d<\log k\end{subarray}}\varphi(d)\leq\sum_{\begin{subarray}{c}d\mid k\\ d<\log k\end{subarray}}d\leq\sum_{\begin{subarray}{c}d\mid k\\ d<\log k\end{subarray}}\log k\leq(\log k)^{2}<\frac{k}{2}.$
The last inequality holds for all sufficiently large $k$, which concludes the proof. ∎
Lemma 62.
Let $(a_{n})_{n\geq 1}$ be a sequence of positive real numbers such that $\inf_{1\leq m\neq n\leq N}|a_{m}-a_{n}|\geq 1$, and $(\eta(N))_{N\geq 1}$ be a sequence of non-negative reals such that $\limsup_{k\to\infty}\eta(k)>0$. Then, the inequality $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$ has infinitely many solutions for almost all $\alpha\in[0,1]$.
Proof.
Let $\limsup_{N\to\infty}\eta(N)=u^{*}>0$. Then, there exist infinitely many $N$ such that $\eta(N)>\frac{u^{*}}{2}$.
Let $A=\displaystyle\bigcup_{N\geq 1}\lfloor(A_{N}-A_{N})^{+}\rfloor$. The spacing condition and Lemma 37 ensure that
(174) $\#A\geq\#\lfloor(A_{N+1}-A_{N+1})^{+}\rfloor=H_{N+1}\geq\frac{N}{2}$
for each $N\in\mathbb{N}$. This shows that $A$ is an infinite set of positive integers.
Therefore, $A$ can be enumerated as $\{Z_{1},Z_{2},\dots\}$, ensuring that for each $N$,
(175) $\{Z_{n}:1\leq n\leq H_{N}\}=\lfloor(A_{N}-A_{N})^{+}\rfloor.$
Since $(Z_{n})_{n\geq 1}$ is a sequence of distinct integers, Weyl’s equidistribution theorem (see Lemma 58) implies that for almost every $\alpha$, the sequence $(\alpha Z_{n})_{n\geq 1}$ is uniformly distributed mod $1$. Hence, for any $\delta>0$,
(176) $\lim_{N\to\infty}\frac{1}{N}\#\left\{k\leq N:\{\alpha Z_{k}\}\in[0,\delta]\right\}=\delta.$
In particular, taking $\delta=\frac{u^{*}}{2}$, for almost every $\alpha$, there exist infinitely many $k$ such that $\{\alpha Z_{k}\}\leq\frac{u^{*}}{2}$, that is, $\|\alpha Z_{k}\|\leq\frac{u^{*}}{2}$.
Since $\eta(N)>\frac{u^{*}}{2}$ for infinitely many $N$, there exist sequences of distinct integers $(k(t))_{t\geq 1}$ and $(N(s))_{s\geq 1}$ such that
(177) $\|\alpha Z_{k(t)}\|\leq\frac{u^{*}}{2}<\eta(N(s))$
for almost every $\alpha$. Since $(N(s))_{s\geq 1}$ is unbounded, there is a sequence of distinct positive integers $(r(t))_{t\geq 1}$ such that $N(r(t))>2k(t)+1$. By Lemma 37, $H_{N(r(t))}\geq\frac{N(r(t))-1}{2}>k(t)$. Therefore, for each $t\in\mathbb{N}$,
(178) $\delta^{\alpha}_{\min}(N(r(t)))=\min_{1\leq m\neq n\leq N(r(t))}\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|=\min_{l\leq H_{N(r(t))}}\|\alpha Z_{l}\|\leq\|\alpha Z_{k(t)}\|\leq\frac{u^{*}}{2}<\eta(N(r(t))).$
Since $(r(t))_{t\geq 1}$ and $(N(s))_{s\geq 1}$ are sequences of distinct integers, we have $(N(r(t))_{t\geq 1}$ is a sequence of distinct integers. Therefore, for almost every $\alpha$, there are infinitely many $N$ such that
(179) $\delta^{\alpha}_{\min}(N)\leq\eta(N).$
Remark 63.
Since a sequence of distinct integers has an increasing subsequence, by passing to one such subsequence $(Z_{l(n)})_{n\geq 1}$ of $(Z_{n})_{n\geq 1}$ and applying Lemma 57, we find that $(\alpha Z_{l(n)})_{n\geq 1}$ is uniformly distributed for almost all $\alpha\in[0,1)$. Hence, Lemma 57 can be used in place of Lemma 58 in the proof of Lemma 62, as noted in [ABM].
Proof of Theorem 13
We break the proof in two cases, depending on whether $\limsup_{k\rightarrow\infty}\eta(k)>0$ or equals 0, where $(\eta(k))_{k\geq 1}$ is the given sequence of non-negative reals.
Case 1: Let $\limsup_{k\to\infty}\eta(k)=u^{*}>0$. We will show that the series (8) diverges and that $\lambda(\mathcal{A})=1$. By Lemma 62, inequality $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$ has infinitely many solutions for almost all $\alpha\in[0,1]$. Consequently, $\lambda(\mathcal{A})=1$.
We now show that the series (8) is indeed divergent in this case. Let $A=\displaystyle\bigcup_{N\geq 1}\lfloor\left(A_{N}-A_{N}\right)^{+}\rfloor$ and fix a number $k\in A$. Then we have $\mathcal{N}(k)<\infty$, where $\mathcal{N}(k)$ is defined in Section 2.1. An application of Lemma 61 yields
(180) $\sum_{\begin{subarray}{c}d\mid k\\ d\geq\log k\end{subarray}}\varphi(d)\sup_{b\geq 1}\left\{\frac{\sup_{\ell\geq\mathcal{N}(bd)}\eta(\ell)}{bd}\right\}$ $\geq\sum_{\begin{subarray}{c}d\mid k\\ d\geq\log k\end{subarray}}\varphi(d)\frac{\sup_{\ell\geq\mathcal{N}(k)}\eta(\ell)}{k}$
(181) $=\frac{\sup_{\ell\geq\mathcal{N}(k)}\eta(\ell)}{k}\left(\sum_{d\mid k}\varphi(d)-\sum_{\begin{subarray}{c}d\mid k\\ d<\log k\end{subarray}}\varphi(d)\right)$
(182) $\geq\frac{\sup_{\ell\geq\mathcal{N}(k)}\eta(\ell)}{2}\geq\frac{u^{*}}{2}>0,$
for sufficiently large $k\in A$, where $\mathcal{N}(k)<\infty$ ensures that the supremum is finite (condition $\mathcal{N}(k)<\infty$ ensures that we are not taking a supremum over an empty set, in which case it would be 0, by convention). Since $A$ is an infinite set, we use (180) and Corollary 60 to deduce that (8) is indeed divergent in this case.
Case 2: Let $\limsup_{k\to\infty}\eta(k)=0$. Since $(\eta(k))_{k\geq 1}$ is a sequence of non-negative reals, we have $\eta(N)\rightarrow 0$. We set
[…]
For every $k\in A$, we set
(183) $S_{k}=[0,1]\cap\left(\bigcup_{0\leq a\leq k}\left(\frac{a}{k}-\frac{\psi(k)}{k},\frac{a}{k}+\frac{\psi(k)}{k}\right)\right),$
and set $S_{k}=\emptyset$ for $k\notin A$. Let
$\mathcal{A}=\{\alpha\in[0,1]:\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)\text{ for infinitely many }N\}$
be defined as in the statement of the theorem, and
$\mathcal{B}=\{\alpha\in[0,1]:\alpha\in S_{k}\text{ for infinitely many }k\}.$
To complete the proof of Theorem 13, we will show that $\mathcal{A}=\mathcal{B}$. By Lemma 56, we have $\lambda(\mathcal{B})=0$ or $\lambda(\mathcal{B})=1$, according to whether the series
(184) $\sum_{k=1}^{\infty}\varphi(k)\sup_{b\geq 1}\left\{\frac{\psi(bk)}{bk}\right\}=\sum_{k=1}^{\infty}\varphi(k)\sup_{b\geq 1}\left\{\frac{\sup_{\ell\geq\mathcal{N}(bk)}\eta(\ell)}{bk}\right\}$
converges or diverges, respectively. Note that $\mathcal{N}(k)=\infty$ for $k\notin A$, and the supremum over the empty set is zero by convention. Since $\eta(\ell)\geq 0$, both expressions
(185) $\displaystyle\sup_{b\geq 1}\left\{\frac{\psi(bk)}{bk}\right\}\quad\text{and}\quad\displaystyle\sup_{b\geq 1}\left\{\frac{\sup_{\ell\geq\mathcal{N}(bk)}\eta(\ell)}{bk}\right\}$
coincide with $\sup_{\begin{subarray}{c}b\geq 1\\ bk\in A\end{subarray}}\left(\frac{\sup_{\ell\geq\mathcal{N}(bk)}\eta(\ell)}{bk}\right)$ for all $k\in\mathbb{N}$. Therefore, the two series in (184) are equal.
To show $\mathcal{A}=\mathcal{B}$, first assume that $\alpha\in\mathcal{A}$. Then, there are infinitely many values of $N,m,n$ with $1\leq m\neq n\leq N$ such that $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$. Set $k=\lfloor|a_{m}-a_{n}|\rfloor$. Since $m,n\leq N$, we have $\mathcal{N}(k)\leq N$. Thus, we have $\psi(k)=\sup_{\ell\geq\mathcal{N}(k)}\eta(\ell)\geq\eta(N)$. Consequently, there are infinitely many values of $k,N$ such that $\|k\alpha\|\leq\eta(N)$, that is, $\|k\alpha\|\leq\psi(k)$, and thus $\alpha\in S_{k}$. Since $\eta(N)\rightarrow 0$, a particular difference $k=\lfloor|a_{m}-a_{n}|\rfloor$ can generate only finitely many values $N$ (together with $m,n\leq N)$ such that $\|\lfloor|a_{m}-a_{n}|\rfloor\alpha\|\leq\eta(N)$. Since each $k$ corresponds to finitely many $N$, and there are infinitely many values of $k,N$ such that $\|k\alpha\|\leq\eta(N)$, the choices for $k$ satisfying $\|k\alpha\|\leq\eta(N)$ must be infinite. In other words, there are infinitely many distinct values of $k$ such that $\alpha\in S_{k}$, and therefore $\mathcal{A}\subseteq\mathcal{B}$.
We now assume that $\alpha\notin\mathcal{A}$. Then, there are only finitely many $N$ such that $\lfloor\delta_{\min}^{\alpha}\rfloor(N)\leq\eta(N)$, that is, there are only finitely many $N$ and $m,n$ with $1\leq m\neq n\leq N$ such that $\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|\leq\eta(N)$. This implies that there are only finitely many $k,N\in\mathbb{N}$ such that $k\in\lfloor(A_{N}-A_{N})^{+}\rfloor$ and $\|k\alpha\|\leq\eta(N)$. Since $\eta(N)\rightarrow 0$, there are only finitely many $k$, and for each such $k$, there are only finitely many $\ell\geq\mathcal{N}(k)$ satisfying $\|k\alpha\|\leq\eta(l)$. Therefore, there are only finitely many $k$ for which $\|k\alpha\|\leq\sup_{l\geq\mathcal{N}(k)}\eta(l)$, that is, $\|k\alpha\|\leq\psi(k)$. Thus, there are only finitely many $k$ such that $\alpha\in S_{k}$, which shows that $\alpha\notin\mathcal{B}$.
Thus, we have shown that $\mathcal{A}=\mathcal{B}$. The convergence or divergence of (8) provides a zero-one law for $\lambda(\mathcal{B})$, and hence also for $\lambda(\mathcal{A})$, as stated in Theorem 13.
Acknowledgements
The author is grateful to Christoph Aistleitner and Sudhir Pujahari for thoughtful inputs and correspondence. The author is supported by the Department of Atomic Energy (DAE project number RIN-4001).